### v1

# Exact and efficient basis pursuit denoising via differential inclusions and a selection principle

Abstract
Basis pursuit denoising (BPDN) is a cornerstone of compressive sensing, statistics and machine learning. While various algorithms for BPDN have been proposed, they invariably suffer from drawbacks and must either favor efficiency at the expense of accuracy or vice versa. As such, state-of-the-art algorithms remain ineffective for high-dimensional applications that require accurate solutions within a reasonable amount of computational time. In this work, we address this issue and propose an exact and efficient algorithm for BPDN using differential inclusions. Specifically, we prove that a selection principle from the theory of differential inclusions turns the dual problem of BPDN into calculating the trajectory of an integrable projected dynamical system, that is, whose trajectory and asymptotic limit can be computed exactly. Our analysis naturally yields an exact algorithm, numerically up to machine precision, that is amenable to computing regularization paths and very fast. Numerical experiments confirm that our algorithm outperforms the state-of-the-art algorithms in both accuracy and efficiency. Moreover, we show that the global continuation of solutions (in terms of the hyperparameter and data) of the projected dynamical system yields a rigorous homotopy algorithm for BPDN, as well as a novel greedy algorithm for computing feasible solutions to basis pursuit in strongly polynomial time. Beyond this work, we expect that our results and analysis can be adapted to compute exact or approximate solutions to a broader class of polyhedral-constrained optimization problems.
Keywords: Basis pursuit denoising, Lasso, $\ell_{1}$-regularized least-squares, compressive sensing, statistics, machine learning, inverse problems, optimization, convex analysis, differential inclusions, projected dynamical systems, variational inequalities, exact solution, algorithms.
Mathematics Subject Classification: 90C25, 65K05, 37N40, 46N10, 34A60, 62J07
1 Introduction
Basis pursuit denoising (BPDN) is an important mathematical optimization problem relevant for high-dimensional applications in, e.g., compressive sensing, statistics, machine learning and inverse problems [17, 61, 48, 69, 58]. Also known as the lasso problem or the $\ell_{1}$-regularized least-squares problem, BPDN refers to the $\ell_{1}$-regularized least-squares problem
$\min_{\boldsymbol{x}\in\mathbb{R}^{n}}\left\{\left\|{\boldsymbol{x}}\right\|_{1}+\frac{1}{2t}\left\|{\bm{A}\boldsymbol{x}-\boldsymbol{b}}\right\|_{2}^{2}\right\},$ (BPDN)
where $\bm{A}$ is a real $m\times n$ matrix with $m\leqslant n$, $\boldsymbol{b}\in\mathbb{R}^{m}$ is the observed data, and $t\geqslant 0$ is a hyperparameter controlling the trade-off between sparsity and data fidelity. The limit $t\to 0$ yields the basis pursuit (BP) problem
$\min_{\boldsymbol{x}\in\mathbb{R}^{n}}\left\|{\boldsymbol{x}}\right\|_{1}\quad\text{subject to}\quad\bm{A}\boldsymbol{x}=\boldsymbol{b}.$ (BP)
The interpretation of $\bm{A}$ and $\boldsymbol{b}$ depends on the application, while the hyperparameter $t$ is either preselected or estimated with data-driven methods such as cross-validation.
Estimating the appropriate hyperparameter can prove challenging in practice. The standard approach is to construct a regularization path. To do so, one first selects a sequence $\{t^{(k)}\}_{k=0}^{K}$, one then computes the corresponding solutions $\{\boldsymbol{x}^{s}(t^{(k)},\boldsymbol{b})\}_{k=0}^{K}$ to (BPDN), and one finally chooses the hyperparameter that gives the preferred solution. Different methods differ in how they select the hyperparameters and compute the solutions, but in any case, constructing a regularization path entails solving (BPDN) accurately for many hyperparameters. This approach can therefore become time-consuming and computationally intensive when the dimensions $m$ and $n$ are high. This issue has stimulated significant research to develop efficient and accurate algorithms for solving (BPDN) in high dimensions [50, 49, 25, 18, 29, 22, 6, 16, 11].
While many algorithms have been proposed for BPDN, they invariably suffer from drawbacks and must either favor efficiency at the expense of accuracy or vice versa. As such, the state-of-the-art algorithms remain ineffective for high-dimensional applications where accurate solutions are required in a reasonable amount of computational time. In this work, we address this issue and propose an exact and efficient algorithm for BPDN using differential inclusions. Our detailed analysis and results yield Algorithm 1 below. It computes an exact pair of primal and dual solutions to BPDN, numerically up to machine precision, is amenable to computing regularization paths, and is very fast.
Input : A matrix $\bm{A}\in\mathbb{R}^{m\times n}$, a vector $\boldsymbol{b}\in\mathrm{Im}({\bm{A}})\setminus{\{\boldsymbol{0}\}}$ and a number $t\geqslant 0$.
Output : A pair of primal and dual solutions $(\boldsymbol{x}^{s}(t,\boldsymbol{b}),\boldsymbol{p}^{s}(t,\boldsymbol{b}))\in\mathbb{R}^{n}\times\mathbb{R}^{m}$ to basis pursuit denoising.
1 Set $\boldsymbol{p}^{(0)}\in\mathbb{R}^{m}$ such that $\left\|{\bm{A}^{\top}\boldsymbol{p}^{(0)}}\right\|_{\infty}=1$;
2 for $k=1$ until convergence do
3       Compute $\mathcal{E}^{(k-1)}=\{j\in\{1,\dots,n\}:|\langle-\bm{A}^{\top}\boldsymbol{p}^{(k-1)},\boldsymbol{e}_{j}\rangle|=1\}$;
4       Compute $\bm{D}^{(k-1)}=\mathrm{diag}\left(\mathrm{sgn}{(-\bm{A}^{\top}\boldsymbol{p}^{(k-1)})}\right)$;
5       Compute $\hat{\boldsymbol{u}}^{(k-1)}\in\operatorname*{arg\,min}_{\begin{subarray}{c}\boldsymbol{u}_{\mathcal{E}^{(k-1)}}\geqslant\boldsymbol{0}\\ \boldsymbol{u}_{{(\mathcal{E}^{(k-1)})^{\mathsf{C}}}}=\boldsymbol{0}\end{subarray}}\left\|{\bm{A}\bm{D}^{(k-1)}\boldsymbol{u}-\boldsymbol{b}-t\boldsymbol{p}^{(k-1)}}\right\|_{2}^{2}$ ;
6       Compute the descent direction $\boldsymbol{d}^{(k-1)}=\bm{A}\bm{D}^{(k-1)}\hat{\boldsymbol{u}}^{(k-1)}-\boldsymbol{b}-t\boldsymbol{p}^{(k-1)}$;
7       Compute the maximal descent time
$\Delta^{(k-1)}=\min_{j\in\{1,\dots,n\}}\left\{\frac{\mathrm{sgn}{\left\langle\bm{D}^{(k-1)}\bm{A}^{\top}\boldsymbol{d}^{(k-1)},\boldsymbol{e}_{j}\right\rangle}-\left\langle\bm{D}^{(k-1)}\bm{A}^{\top}\boldsymbol{p}^{(k-1)},\boldsymbol{e}_{j}\right\rangle)}{\left\langle\bm{D}^{(k-1)}\bm{A}^{\top}\boldsymbol{d}^{(k-1)},\boldsymbol{e}_{j}\right\rangle}\right\}\;$
8      if $t>0$ and $t\Delta^{(k-1)}\geqslant 1$ then
9             Set $\boldsymbol{x}^{s}(t,\boldsymbol{b})=\bm{D}^{(k-1)}\hat{\boldsymbol{u}}^{(k-1)}$;
10             Set $\boldsymbol{p}^{s}(t,\boldsymbol{b})=\boldsymbol{p}^{(k-1)}+\boldsymbol{d}^{(k-1)}/t$;
11             break; // The algorithm has converged
12       else if $t=0$ and $\boldsymbol{d}^{(k-1)}=\boldsymbol{0}$ then
13             Set $\boldsymbol{x}^{s}(t,\boldsymbol{b})=\bm{D}^{(k-1)}\hat{\boldsymbol{u}}^{(k-1)}$;
14             Set $\boldsymbol{p}^{s}(t,\boldsymbol{b})=\boldsymbol{p}^{(k-1)}$;
15             break; // The algorithm has converged
16       Update $\boldsymbol{p}^{(k)}=\boldsymbol{p}^{(k-1)}+\Delta^{(k-1)}\boldsymbol{d}^{(k-1)}$;
17
18 end for
Algorithm 1 Algorithm for computing a pair of primal and dual solutions $(\boldsymbol{x}^{s}(t,\boldsymbol{b}),\boldsymbol{p}^{s}(t,\boldsymbol{b}))$ to basis pursuit denoising.
Contributions of this paper:
(i) In this work, we prove that a selection principle from the theory of differential inclusions turns the dual problem of BPDN into calculating the trajectory of an integrable projected dynamical system, that is, whose trajectory and asymptotic limit can be computed exactly. Our results and extensive analysis culminate in Algorithm 1, an active set algorithm that computes the optimal dual solution to BPDN and recovers an optimal primal solution exactly. As a by-product, our results provide a novel solution method for solving a broad class of projected dynamical systems, which should find relevance to applications outside BPDN.
(ii) We present a detailed perturbation analysis of solutions to BPDN (in terms of the hyperparameter and data) via a global continuation argument of solutions of the projected dynamical system characterizing BPDN. This yields a rigorous generalized homotopy algorithm for BPDN, as well as a novel greedy algorithm for computing feasible solutions to basis pursuit in strongly polynomial time.
(iii) Our numerical experiments show that Algorithm 1 vastly outperforms the state of the arts in accuracy while also achieving the best overall performance, highlighting a key feature of Algorithm 1: it neither compromises accuracy nor computational efficiency.
Related work:
The state-of-the-art algorithms for BPDN are divided roughly in three categories: coordinate descent methods [29, 71, 30, 62, 70, 57, 27, 37], first-order optimization algorithms such as FISTA or the PDHG algorithm [33, 72, 6, 9, 15, 16], and homotopy algorithms [50, 49, 25, 22, 63, 11]. We will focus here only on these algorithms; a partial list of other algorithms include iteratively reweighted least squares algorithms [18, 56], approximate online algorithms [4, 52, 31], Bayesian methods [34, 36, 51, 42, 1], the adaptive inverse scale space method [14], specialized Newton and quasi-Newton methods [73, 40, 66], fixed-point continuation methods [35], and interior point methods [10]. See [75, 7, 43, 68] for surveys and comparisons of different methods and models.
Coordinate descent methods are the state of the arts for high-dimensional BPDN because they are fast. Key to their efficiency are so-called selection rules or heuristics [62, 32, 53] that estimate, a priori, the degree of sparsity of primal solutions. These methods, however, are known to suffer from algorithmic instability [5] and therefore may produce inaccurate and unreliable numerical solutions. First-order optimization algorithms, in contrast, are numerically accurate but less efficient in high dimensions. Finally, homotopy algorithms compute exact solutions paths to (BPDN) but generally require technical assumptions to work, e.g., the uniqueness of the path and the “one-at-a-time condition” [63, 25, 50]. These assumptions are difficult to verify and may not hold in practice. For example, the LARS algorithm fails to converge on simple examples [11, Proposition 4.1]. Moreover, all homotopy algorithms have exponential worst-case complexity [45], and while in practice they often converge somewhat fast, they are generally less efficient than coordinate descent methods.
Organization of this paper:
We review in Section 2 the existence of solutions and optimality conditions to BPDN and introduce a characterization in terms of differential inclusions. In Section 3, we present the minimal selection principle, use it to cast the dual problem of BPDN into the equivalent problem of computing the trajectory of a projected dynamical system, and then characterize various properties of the system, including its evolution rule. Then, we use this evolution rule in Section 4 to prove that the projected dynamical system can be integrated and therefore its trajectory and asymptotic limit calculated explicitly. This gives the optimal solution to the dual problem of BPDN when $t>0$ (an optimal solution when $t=0$) and recover an optimal primal solution exactly. We present in Section 5 a detailed perturbation analysis of the trajectory of the projected dynamical system, equivalently the solutions to BPDN, in terms of the hyperparameter $t$ and data $\boldsymbol{b}$. We prove that this global continuation yields a generalized homotopy algorithm for BPDN. Furthermore, we use the global continuation in Section 6 to devise a greedy algorithm for computing feasible primal and dual solutions to BPDN in strongly polynomial time. We present numerical experiments in Section 7 to compare our algorithms to some state-of-the-art algorithms for BPDN, and finally we discuss the broader implications of our work in Section 8.
2 Setup
Unless stated otherwise, we assume $\boldsymbol{b}\in\text{Im}(\bm{A})\setminus\{\boldsymbol{0}\}$ and $\mathrm{rank}({\bm{A}})=m$. The case $\boldsymbol{b}=\boldsymbol{0}$ is uninteresting and we can assume the latter because, otherwise, at least one row of the matrix $\bm{A}$ is linearly dependent on the others and can therefore be discarded.
We wish to note that some analyses and proofs in this paper are fairly involved and use concepts that may be unfamiliar to the reader (e.g., advanced notions of convex analysis, functional analysis, and dynamical systems theory). We refer the reader to Appendix A for more detailed mathematical background.
2.1 Existence of solutions and optimality conditions
We review here the existence of solutions and optimality conditions to BPDN using classic results from convex analysis, summarized as Theorem A.3 in Appendix A. Following the notation of Theorem A.3, we set
$f_{1}(\cdot)=\left\|{\cdot}\right\|_{1}\quad\text{and}\quad f_{2}(\cdot)=\begin{cases}\frac{1}{2t}\left\|{\cdot-\boldsymbol{b}}\right\|_{2}^{2}\,&\text{if }t>0\\ \chi_{\{b\}}(\cdot)\,&\text{if }t=0\end{cases}$
where $\chi_{\boldsymbol{b}}(\cdot)$ denotes the characteristic function of the singleton set $\{\boldsymbol{b}\}$. Since $\mathrm{dom~{}}f_{1}=\mathbb{R}^{n}$, $\boldsymbol{b}\in\mathrm{ri}\,\mathrm{dom~{}}f_{2}$ and $\boldsymbol{b}\in\mathrm{Im}({\bm{A}})$, there exists $\boldsymbol{x}\in\mathrm{ri}\,\mathrm{dom~{}}f_{1}$ such that $\bm{A}\boldsymbol{x}\in\mathrm{ri}\,\mathrm{dom~{}}f_{2}$. Moreover, the function $\boldsymbol{x}\mapsto f_{1}(\boldsymbol{x})+f_{2}(\bm{A}\boldsymbol{x})$ is coercive because $f_{1}$ is coercive and $f_{2}$ is nonnegative. Using Remark A.1, we conclude that all conditions of Theorem A.3 are satisfied. As a consequence, (BPDN) and (BP) both have at least one solution. The problem
$-\inf_{\boldsymbol{p}\in\mathbb{R}^{m}}\left\{\frac{t}{2}\left\|{\boldsymbol{p}}\right\|_{2}^{2}+\langle\boldsymbol{p},\boldsymbol{b}\rangle+\chi_{\mathrm{B}_{\infty}}(-\bm{A}^{\top}\boldsymbol{p})\right\}$ (dBPDN)
is the dual of BPDN (dBPDN) and the problem
$-\inf_{\boldsymbol{p}\in\mathbb{R}^{m}}\left\{\langle\boldsymbol{p},\boldsymbol{b}\rangle+\chi_{\mathrm{B}_{\infty}}(-\bm{A}^{\top}\boldsymbol{p})\right\}$ (dBP)
is the dual of BP (dBP). Here, $\chi_{B_{\infty}}(\cdot)$ denotes the characteristic function of the unit $\ell_{\infty}$-ball:
$\chi_{\mathrm{B}_{\infty}}(\boldsymbol{s})=\begin{cases}0\,&\,\text{if }|\langle\boldsymbol{s},\boldsymbol{e}_{j}\rangle|\leqslant 1\,\text{for every}\,j\in\{1,\dots,n\},\\ +\infty\,&\,\text{otherwise}.\end{cases}$ (1)
Problem (dBPDN) has a unique solution for every $t>0$ due to strong convexity and problem (dBP) has at least one solution. Moreover, (dBPDN) and (dBP) are equal in value to their respective primal problems (BPDN) and (BP). Finally, letting $(\boldsymbol{x}^{s}(t,\boldsymbol{b}),\boldsymbol{p}^{s}(t,\boldsymbol{b}))$ denote any pair of solutions to basis pursuit (denoising) and its dual, we have the following set of equivalent first-order optimality conditions:
$-t\boldsymbol{p}^{s}(t,\boldsymbol{b})$ $=\boldsymbol{b}-\bm{A}\boldsymbol{x}^{s}(t,\boldsymbol{b})\quad\text{and}\ -\bm{A}^{\top}\boldsymbol{p}^{s}(t,\boldsymbol{b})\in\partial\left\|{\cdot}\right\|_{1}(\boldsymbol{x}^{s}(t,\boldsymbol{b})),$ (2a)
$-t\boldsymbol{p}^{s}(t,\boldsymbol{b})$ $\in\boldsymbol{b}-\bm{A}\partial\chi_{\mathrm{B}_{\infty}}(-\bm{A}^{\top}\boldsymbol{p}^{s}(t,\boldsymbol{b})),$ (2b)
where $\lim_{t\to 0}t\boldsymbol{p}^{s}(t,\boldsymbol{b})=0$, with $\bm{A}\boldsymbol{x}^{s}(0)=\boldsymbol{b}$ and $-\bm{A}^{\top}\boldsymbol{p}^{s}(0)\in\partial\left\|{\cdot}\right\|_{1}(\boldsymbol{x}^{s}(0))$.
2.2 Structure of the optimality conditions
The optimality conditions in (2a) on the right identify the components of $-\bm{A}^{\top}\boldsymbol{p}^{s}(t,\boldsymbol{b})$ that achieve maximum absolute deviation:
$\langle-\bm{A}^{\top}\boldsymbol{p}^{s}(t,\boldsymbol{b}),\boldsymbol{e}_{j}\rangle=\begin{cases}1\,&\,\text{if }x_{j}^{s}(t,\boldsymbol{b})>0\\ -1\,&\,\text{if }x_{j}^{s}(t,\boldsymbol{b})<0\\ [-1,1]\,&\,\text{if }x_{j}^{s}(t,\boldsymbol{b})=0.\end{cases}$ (3)
In particular, $|\langle-\bm{A}^{\top}\boldsymbol{p}^{s}(t,\boldsymbol{b}),\boldsymbol{e}_{j}\rangle|<1\implies x_{j}^{s}(t,\boldsymbol{b})=0$. It will therefore be useful to identify for any $\boldsymbol{p}\in\mathbb{R}^{m}$ the set of indices $j\in\{1,\dots,n\}$ for which $|\langle-\bm{A}^{\top}\boldsymbol{p},\boldsymbol{e}_{j}\rangle|=1$. This set is called the equicorrelation set at $\boldsymbol{p}$ and we will denote it by $\mathcal{E}(\boldsymbol{p})$:
$\mathcal{E}(\boldsymbol{p})\coloneqq\left\{j\in\{1,\dots,n\}:|\langle-\bm{A}^{\top}\boldsymbol{p},\boldsymbol{e}_{j}\rangle|=1\right\}.$ (4)
In addition, we will keep track of the signs and define the $n\times n$ diagonal matrix of signs
$\bm{D}(\boldsymbol{p})=\mathrm{diag}\left(\mathrm{sgn}{(-\bm{A}^{\top}\boldsymbol{p})}\right).$ (5)
The equicorrelation set identifies the active constraints of the cone (2b). To see this, observe the $\ell_{\infty}$ unit ball is a closed convex polyhedron, and so Proposition (A.1) implies
$\boldsymbol{b}-\bm{A}\partial\chi_{\mathrm{B}_{\infty}}(-\bm{A}^{\top}\boldsymbol{p})=\left\{\boldsymbol{b}-\sum_{j\in\mathcal{E}(\boldsymbol{p})}u_{j}\bm{A}\bm{D}(\boldsymbol{p})\boldsymbol{e}_{j}:u_{j}\geqslant 0\right\}$
for all $\boldsymbol{p}\in\mathrm{dom~{}}V(\cdot;t,\boldsymbol{b})$. The optimality conditions (2b) therefore read
$-t\boldsymbol{p}^{s}(t,\boldsymbol{b})\in\left\{\boldsymbol{b}-\sum_{j\in\mathcal{E}(\boldsymbol{p}^{s}(t,\boldsymbol{b}))}u_{j}\bm{A}\bm{D}(\boldsymbol{p}^{s}(t,\boldsymbol{b}))\boldsymbol{e}_{j}:u_{j}\geqslant 0\right\}.$ (6)
2.3 Characterization via differential inclusions
We now present a characterization of solutions to the dual problem (dBPDN) in terms of an initial value problem involving differential inclusions. Let $V\colon\mathbb{R}^{m}\times[0,+\infty)\times\mathbb{R}^{m}\to\mathbb{R}\cup\{+\infty\}$ denote the objective function in (dBPDN), parameterized in terms of $t$ and $\boldsymbol{b}$:
$V(\boldsymbol{p};t,\boldsymbol{b})=\frac{t}{2}\left\|{\boldsymbol{p}}\right\|_{2}^{2}+\langle\boldsymbol{p},\boldsymbol{b}\rangle+\chi_{\mathrm{B}_{\infty}}(-\bm{A}^{\top}\boldsymbol{p}).$ (7)
Its subdifferential with respect to $\boldsymbol{p}$ is the convex cone
$\partial_{\boldsymbol{p}}V(\boldsymbol{p};t,\boldsymbol{b})=\left\{t\boldsymbol{p}+\boldsymbol{b}-\sum_{j\in\mathcal{E}(\boldsymbol{p})}u_{j}\bm{A}\bm{D}(\boldsymbol{p})\boldsymbol{e}_{j}:u_{j}\geqslant 0\right\}.$ (8)
(See Proposition A.1 for details.) We suggest to compute a solution to the dual problem using a nonsmooth generalization of gradient descent on $\boldsymbol{p}\mapsto V(\boldsymbol{p};t,\boldsymbol{b})$:
$\frac{d\boldsymbol{p}}{d\tau}(\tau)\in-\partial_{\boldsymbol{p}}V(\boldsymbol{p}(\tau);t,\boldsymbol{b}),\qquad\boldsymbol{p}(0)=\boldsymbol{p}_{0}\in\mathrm{dom~{}}\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b}).$ (9)
Differential inclusions generalize the concept of ordinary differential equations to multi-valued maximal monotone mappings [3]. These mappings include, as a special case, the subdifferentials of proper, lower semicontinuous and convex functions (see Definition A.12). Their corresponding differential inclusions are sometimes called gradient inclusions [13, Section 1.3.7, page 10] because they generalize the classical concept of gradient systems [^1] Note that $V(\cdot;t,\boldsymbol{b})$ is differentiable at $\boldsymbol{p}(\tau)\in\mathrm{dom~{}}\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b})\iff\partial_{\boldsymbol{p}}V(\boldsymbol{p}(\tau);t,\boldsymbol{b})=\{\nabla_{\boldsymbol{p}}V(\boldsymbol{p}(\tau);t,\boldsymbol{b})\}$, in which case $\frac{d\boldsymbol{p}}{d\tau}(\tau)=-\nabla_{\boldsymbol{p}}V(\boldsymbol{p}(\tau);t,\boldsymbol{b})$.. Indeed, similarly to how trajectories of gradient systems converge to their critical points (if any exist), the trajectories of gradient inclusions converge to their critical points (if any exist). This fact follows from a variational principle called the minimal selection principle.
3 The minimal selection principle
The minimal selection principle stipulates that solutions to gradient inclusions exist, are unique, and evolve in the steepest descent direction [3, Chapter 3]. As the minimal selection principle is central to this work, we state it below in full:
Theorem 3.1 (Existence and uniqueness of solutions to differential inclusions with maximal monotone mappings).
Let $g\colon\mathbb{R}^{m}\to\mathbb{R}\cup\{+\infty\}$ be a proper, lower semicontinuous, and convex function and let $\boldsymbol{p}_{0}\in\mathrm{dom~{}}\partial g$. Consider the differential inclusions
$\frac{d\boldsymbol{p}}{d\tau}(\tau)\in-\partial g(\boldsymbol{p}(\tau)),\qquad\boldsymbol{p}(0)=\boldsymbol{p}_{0}.$ (10)
Then, there exists a unique solution $\boldsymbol{p}\colon[0,+\infty)\to\mathrm{dom~{}}\partial g$ satisfying (10). Moreover:
  • (i)
    For every $\tau\geqslant 0$, we have $\frac{d\boldsymbol{p}}{d\tau}(\tau)=\lim_{\begin{subarray}{c}h\to 0\\ h>0\end{subarray}}\frac{\boldsymbol{p}(\tau+h)-\boldsymbol{p}(\tau)}{h}$. Moreover, $\tau\mapsto\frac{d\boldsymbol{p}}{d\tau}(\tau)$ is right-continuous and $\tau\mapsto\left\|{\frac{d\boldsymbol{p}}{d\tau}(\tau)}\right\|_{2}$ is nonincreasing.
  • (ii)
    (The minimal selection principle.) The function $\boldsymbol{p}(\cdot)$ satisfies the initial value problem
    $\frac{d\boldsymbol{p}}{d\tau}(\tau)=-\mathrm{proj}_{\partial g(\boldsymbol{p}(\tau))}(\boldsymbol{0}),\qquad\boldsymbol{p}(0)=\boldsymbol{p}_{0}$ (11)
    at $\tau=0$ and almost everywhere on $(0,+\infty)$.
  • (iii)
    (Asymptotic limit.) If $g$ achieves its minimum at some point, then $\boldsymbol{p}(\cdot)$ converges to such a point:
    $\lim_{\tau\to\infty}g(\boldsymbol{p}(\tau))=\min_{\boldsymbol{p}\in\mathbb{R}^{m}}g(\boldsymbol{p})\quad\text{and}\quad\lim_{\tau\to\infty}\boldsymbol{p}(\tau)\in\operatorname*{arg\,min}_{\boldsymbol{p}\in\mathbb{R}^{m}}g(\boldsymbol{p}).$
Proof.
See [3, Theorem 1, page 147 and Theorem 1, page 159] for the proofs of (i)–(iii). All proofs rely on the fact that the subdifferential $\partial g$ is a maximal monotone mapping (see Definition A.12) and build on most of the material presented in [3, Chapters 1-3]. ∎
Remark 3.1.
In the language of dynamical systems theory, the initial value problem (11) is called a projected dynamical system [74]. Such systems arise naturally in the theory of variational inequalities [2, Chapter 17], as was first noted in [23] and later proved in [24].
3.1 The minimal selection principle for basis pursuit denoising
In the context of this work, the minimal selection principle implies that the gradient inclusions (9) has a unique, global solution satisfying the initial value problem
$\frac{d\boldsymbol{p}}{d\tau}(\tau)=-\mathrm{proj}_{\partial_{\boldsymbol{p}}V(\boldsymbol{p}(\tau);t,\boldsymbol{b})}(\boldsymbol{0}),\qquad\boldsymbol{p}(0)\in\mathrm{dom~{}}\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b})$ (12)
at $\tau=0$ and almost everywhere on $(0,+\infty)$, where
$\mathrm{proj}_{\partial_{\boldsymbol{p}}V(\boldsymbol{p}(\tau);t,\boldsymbol{b})}(\boldsymbol{0})=\operatorname*{arg\,min}_{\boldsymbol{s}\in\mathbb{R}^{m}}\left\|{\boldsymbol{s}}\right\|_{2}^{2}\ \text{such that}\ \boldsymbol{s}\in\partial_{\boldsymbol{p}}V(\boldsymbol{p}(\tau);t,\boldsymbol{b}).$
Moreover, the solution converges asymptotically to the unique solution of (dBPDN) when $t>0$ (a solution of (dBP) when $t=0$). We call the projected dynamical system (12) the slow system and its solution the slow solution. In addition, we write
$\boldsymbol{d}(\boldsymbol{p}(\tau);t,\boldsymbol{b})\coloneqq-\mathrm{proj}_{\partial_{\boldsymbol{p}}V(\boldsymbol{p}(\tau);t,\boldsymbol{b})}(\boldsymbol{0})$ (13)
to denote the direction of change of the slow system at $\boldsymbol{p}(\tau)$.
The minimal selection principle suggests we can compute the solution to (dBPDN) (a solution to (dBP) when $t=0$) by calculating the trajectory of the slow system (12) and its asymptotic limit. To do this, we must compute the minimal selection $\mathrm{proj}_{\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0};t,\boldsymbol{b})}(\boldsymbol{0})$ for any $\boldsymbol{p}_{0}\in\mathrm{dom~{}}\partial_{\boldsymbol{p}}V(\cdot,t,\boldsymbol{b})$. We turn to this problem next.
3.2 Computing the minimal selection
Here, we describe how the minimal selection $\mathrm{proj}_{\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0};t,\boldsymbol{b})}(\boldsymbol{0})\equiv-\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ of the slow system (12) is computed from a cone projection problem, equivalently a nonnegative least-squares (NNLS) problem. To state this succinctly, we will use submatrix notation: Given $\boldsymbol{p}\in\mathrm{dom~{}}V$, we denote by $\bm{A}_{\mathcal{E}(\boldsymbol{p})}$ the $m\times|\mathcal{E}(\boldsymbol{p})|$ the submatrix of $\bm{A}$ with columns indexed by $\mathcal{E}(\boldsymbol{p})$, we write $\bm{A}^{\top}_{\mathcal{E}(\boldsymbol{p})}$ to denote its transpose, and we denote by $\bm{D}_{\mathcal{E}(\boldsymbol{p})}$ the $|\mathcal{E}(\boldsymbol{p})|\times|\mathcal{E}(\boldsymbol{p})|$ submatrix of signs $\bm{D}(\boldsymbol{p})$ with rows and columns indexed by $\mathcal{E}(\boldsymbol{p})$. Finally, we denote by $\boldsymbol{u}_{\mathcal{E}(\boldsymbol{p})}$ the subvector of $\boldsymbol{u}\in\mathbb{R}^{n}$ indexed by $\mathcal{E}(\boldsymbol{p})$. With the notation set, we have the following:
Lemma 3.1.
Let $t\geqslant 0$ and $\boldsymbol{p}_{0}\in\mathrm{dom~{}}\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b})$. The direction of change $\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ of the slow system (12) is the unique solution to the cone projection problem
$\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})=\operatorname*{arg\,min}_{\boldsymbol{d}\in\mathbb{R}^{m}}\left\|{\boldsymbol{d}+(\boldsymbol{b}+t\boldsymbol{p}_{0})}\right\|_{2}^{2}\ \text{subject to}\ \bm{D}_{\mathcal{E}(\boldsymbol{p}_{0})}\bm{A}^{\top}_{\mathcal{E}(\boldsymbol{p}_{0})}\boldsymbol{d}\geqslant\boldsymbol{0}.$ (14)
It admits the characterization
$\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})=\bm{A}_{\mathcal{E}(\boldsymbol{p}_{0})}\bm{D}_{\mathcal{E}(\boldsymbol{p}_{0})}\hat{\boldsymbol{u}}_{\mathcal{E}(\boldsymbol{p}_{0})}(\boldsymbol{p}_{0};t,\boldsymbol{b})-(\boldsymbol{b}+t\boldsymbol{p}_{0}),$ (15)
where $\hat{\boldsymbol{u}}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ is a solution to the NNLS problem
$\hat{\boldsymbol{u}}(\boldsymbol{p}_{0};t,\boldsymbol{b})\in\operatorname*{arg\,min}_{\boldsymbol{u}\in\mathbb{R}^{n}}\left\|{\bm{A}\bm{D}(\boldsymbol{p}_{0})\boldsymbol{u}-(\boldsymbol{b}+t\boldsymbol{p}_{0})}\right\|_{2}^{2}\ \text{subject to}\ \begin{cases}&\boldsymbol{u}_{\mathcal{E}(\boldsymbol{p}_{0})}\geqslant\boldsymbol{0},\\ &\boldsymbol{u}_{\mathcal{E}^{\mathsf{C}}(\boldsymbol{p}_{0})}=\boldsymbol{0}.\end{cases}$ (16)
Moreover:
$\hat{\boldsymbol{u}}_{j}(\boldsymbol{p}_{0};t,\boldsymbol{b})[\bm{D}(\boldsymbol{p}_{0})\bm{A}^{\top}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})]_{j}=0\ \text{for every }j\in\{1,\dots,n\},$ (17a)
$\left\|{\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})}\right\|_{2}^{2}+\langle\boldsymbol{b}+t\boldsymbol{p}_{0},\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\rangle=0.$ (17b)
Proof.
For every $t\geqslant 0$ and $\boldsymbol{p}_{0}\in\mathrm{dom~{}}\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b})$, the subdifferential $\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0};t,\boldsymbol{b})$ is non-empty, closed and convex. Hence the projection of $\boldsymbol{0}$ onto $\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0};t,\boldsymbol{b})$ exists and is unique (see Definition A.17). The projection is given precisely the NNLS problem (16) or equivalently by its dual problem, the cone projection problem (14).
$\langle\hat{\boldsymbol{u}}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\bm{D}(\boldsymbol{p}_{0})\bm{A}^{\top}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\rangle$ $=\langle\bm{A}\bm{D}(\boldsymbol{p}_{0})\hat{\boldsymbol{u}}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\rangle$
$=\langle\bm{A}\bm{D}(\boldsymbol{p}_{0})\hat{\boldsymbol{u}}(\boldsymbol{p}_{0};t,\boldsymbol{b})-\boldsymbol{b}-t\boldsymbol{p},\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\rangle$
$\qquad\qquad+\langle\boldsymbol{b}+t\boldsymbol{p},\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\rangle$
$=\left\|{\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})}\right\|_{2}^{2}+\langle\boldsymbol{b}+t\boldsymbol{p},\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\rangle,$
where the last line follows by (15). Finally, combine (15) with (17a) to deduce (17b). ∎
Remark 3.2.
The NNLS problem (16) may have more than one optimal solution, but the direction $\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ is always unique. A detailed review of cone projection algorithms suitable for (14) can be found in [20]. Algorithms for NNLS problems include active set algorithms with “finite-time” convergence, such as the Lawson–Hanson algorithm [41] (see also [12, 19]) and its (highly underappreciated) generalization, Meyer’s algorithm [47].
3.3 The minimal selection yields a descent direction
Lemma 3.1 shows that the minimal selection $-\mathrm{proj}_{\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0};t,\boldsymbol{b})}(\boldsymbol{0})$, equivalently the instantaneous direction $\boldsymbol{d}(\boldsymbol{p}(\tau);t,\boldsymbol{b})$ of the slow system (12), is calculated from a cone projection problem. The next proposition shows that this characterization implies $\boldsymbol{d}(\boldsymbol{p}(\tau);t,\boldsymbol{b})$ is a descent direction for $V$ and, crucially, obeys an evolution rule.
Proposition 3.1.
Let $t\geqslant 0$ and $\boldsymbol{p}_{0}\in\mathrm{dom~{}}\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b})$. Then:
  • (i)
    There exists $\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})>0$ such that $\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\in\mathrm{dom~{}}V(\cdot;t,\boldsymbol{b})$ for every $\Delta\in[0,\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})]$, where
    $\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})=\min_{j\in\{1,\dots,n\}}\left\{\frac{\mathrm{sgn}{\left\langle\bm{D}(\boldsymbol{p}_{0})\bm{A}^{\top}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\boldsymbol{e}_{j}\right\rangle}-\left\langle\bm{D}(\boldsymbol{p}_{0})\bm{A}^{\top}\boldsymbol{p}_{0},\boldsymbol{e}_{j}\right\rangle)}{\left\langle\bm{D}(\boldsymbol{p}_{0})\bm{A}^{\top}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\boldsymbol{e}_{j}\right\rangle}\right\}.$ (18)
    Moreover, $\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})=+\infty\iff\bm{A}^{\top}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})=\boldsymbol{0}\iff\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})=\boldsymbol{0}$.
  • (ii)
    (Descent direction) For every $\Delta\in[0,\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})]$, we have
    $V(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b});t,\boldsymbol{b})-V(\boldsymbol{p}_{0};t,\boldsymbol{b})=\Delta\left(t\Delta/2-1\right)\left\|{\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})}\right\|_{2}^{2}.$ (19)
    In particular, if $\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\neq\boldsymbol{0}$, then it is a descent direction of $V(\cdot;t,\boldsymbol{b})$.
  • (iii)
    For every $\Delta\in[0,\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})]$, we have the inclusion
    $-(1-t\Delta)\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\in\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b});t,\boldsymbol{b}).$
  • (iv)
    For every $\Delta\in[0,\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$, we have the inclusions
    $\mathcal{E}(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}))\subset\mathcal{E}(\boldsymbol{p}_{0})$
    and
    $\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b});t,\boldsymbol{b})\subset\{t\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\}+\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0};t,\boldsymbol{b}).$
  • (v)
    (Evolution rule) For every $\Delta\in[0,\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$, the following evolution rule holds:
    $\boldsymbol{d}(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b});t,\boldsymbol{b})=(1-t\Delta)\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}).$ (20)
Proof.
See Appendix B.1 for the proof. ∎
Remark 3.3 (Drift to infinity).
Without the full row rank assumption $\mathrm{rank}({\bm{A}})=m$, there may be some $\boldsymbol{p}_{0}\in\mathrm{dom~{}}\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b})$ for which $\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\neq\boldsymbol{0}$ with $\bm{A}^{\top}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})=\boldsymbol{0}$. In this case, $\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\in\mathrm{dom~{}}\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b})$ for every $\Delta\in[0,+\infty)$, hence $\lim_{\Delta\to+\infty}V(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}))=+\infty$ by (18). This will have implications for the calculation of the trajectory of the slow system (12) when $\mathrm{rank}({\bm{A}})\neq m$.
4 Dynamics of the slow system
4.1 Explicit local solution
The evolution rule (20) describes how the descent direction $\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ evolves locally in $\Delta>0$ along the line $\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})$. This evolution is local because it holds on some possibly finite interval $\Delta\in[0,\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$, with $\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ given by (18). This behavior, a consequence of the minimal selection principle, fundamentally characterizes the dynamics of the slow system (12). It also appears as the main obstacle for computing the slow solution.
Here, we show that this behavior is no obstacle as it appears: Locally, the slow system evolves as that of its non-projected counterpart, in the sense that
$\frac{d\boldsymbol{p}(\tau)}{d\tau}=-\mathrm{proj}_{\partial_{\boldsymbol{p}}V(\boldsymbol{p}(\tau);t,\boldsymbol{b})}(\boldsymbol{0})=-t(\boldsymbol{p}(\tau)-\boldsymbol{p}_{0})+\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\ \text{ for small enough times $\tau>0$.}$
Its local solution can therefore be computed explicitly, as the following Theorem makes precise.
Theorem 4.1.
Let $t\geqslant 0$ and $\boldsymbol{p}_{0}\in\mathrm{dom~{}}\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b})$. The slow system (12) with initial value $\boldsymbol{p}(0)=\boldsymbol{p}_{0}$ coincides with the initial value problem
$\frac{d\boldsymbol{p}}{d\tau}(\tau)=e^{-t\tau}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\quad\boldsymbol{p}(0)=\boldsymbol{p}_{0}$ (21)
on $\tau\in[0,\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$, where
$\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})=\begin{cases}\Delta_{*}(\boldsymbol{p}_{0};0,\boldsymbol{b})&\text{if}\ t=0,\\ -\ln\left(1-t\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})\right)/t&\text{if}\ t>0\ \text{and}\ 1-t\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})>0,\\ +\infty&\text{otherwise}.\end{cases}$ (22)
In particular, the slow solution is given explicitly on $[0,\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$ by
$\boldsymbol{p}(\tau)=\boldsymbol{p}_{0}+f(\tau,t)\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),$ (23)
where
$f(\tau,t)=\begin{cases}\tau&\text{if}\ t=0,\\ \left(1-e^{-t\tau}\right)/t&\text{if}\ t>0.\end{cases}$ (24)
Proof.
We will use the evolution rule (20) to show that the affine system
$\frac{d\boldsymbol{p}_{a}}{d\tau}(\tau)=-t(\boldsymbol{p}_{a}(\tau)-\boldsymbol{p}_{0})+\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\qquad\boldsymbol{p}_{a}(0)=\boldsymbol{p}_{0}.$ (25)
satisfies the slow system (12) on $[0,\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$ and conclude using uniqueness.
First, consider the affine system (25). A short calculation shows that its unique, global solution is given by
$\boldsymbol{p}_{a}(\tau)=\boldsymbol{p}_{0}+f(\tau,t)\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}).$
Substitute the solution in (25) above to find
$\frac{d\boldsymbol{p}_{a}}{d\tau}(\tau)$ $=(1-tf(\tau,t))\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ (26)
$=e^{-t\tau}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}).$
Next, we invoke the evolution rule (20) in Proposition 3.1(v) with $\Delta=f(\tau,t)$ to find
$-\mathrm{proj}_{\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0}+f(\tau,t)\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),t)}(\boldsymbol{0})=(1-tf(\tau,t))\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ (27)
whenever $0\leqslant f(\tau,t)<\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$. Notice $\tau\mapsto f(\tau,t)$ increases monotonically from 00 to $1/t$ (0 to +$\infty$+ ∞) when $t>0$ ($t=0$). Hence if $t=0$, the largest value of $\tau$ for which (27) holds is $\Delta_{*}(\boldsymbol{p}_{0};0,\boldsymbol{b})$. If $t>0$ and $01-t\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})>$ 0, then $(1-e^{-t\tau})/t$ is equal to $\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ at $-\ln\left(1-t\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})\right)/t$. If $t>0$ and $1-t\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))\leqslant 0$, then (27) holds for every $\tau\geqslant 0$. Taken together, we find
$\frac{d\boldsymbol{p}_{a}}{d\tau}(\tau)=-\mathrm{proj}_{\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0}+f(\tau,t)\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),t)}(\boldsymbol{0})\ \text{for every $\tau\in[0,\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$}.$
Uniqueness follows from Theorem 3.1, and hence (21) follows from (26). ∎
4.2 Continuation of the local solution
Theorem 4.1 provides the explicit local solution to the slow system (12) on the interval $[0,\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$. When $\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})<+\infty$, what can we say about the slow solution and its time derivative at $\tau=\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$? The next result shows that the slow solution is continuous at $\tau=\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ but its time derivative is not.
Lemma 4.1.
Let $t\geqslant 0$, $\boldsymbol{p}_{0}\in\mathrm{dom~{}}\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b})$, and suppose $\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})<+\infty$. Then (23) holds at $\tau=\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ but (21) does not:
$\left\|{\boldsymbol{d}(\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}));t,\boldsymbol{b})}\right\|_{2}<e^{-t\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})}\left\|{\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})}\right\|_{2}.$ (28)
Proof.
See Appendix B.2 for the proof. ∎
4.3 Explicit global solution
So far, we have computed the explicit local solution of the slow system (12) on an interval $[0,\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$ which, when $\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ is finite, can be extended to $\tau=\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$. We can apply Theorem 4.1 and Lemma 4.1 again with the initial condition $\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$ to extend the local solution to $[0,\tau_{*}(\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}));t,\boldsymbol{b})]$. This is because $\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))\in\mathrm{dom~{}}\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b})$ and all assumptions of Theorem 4.1 hold.
We now apply this argument repeatedly to compute explicitly the global solution of the slow system. Let $\boldsymbol{p}^{(0)}\in\mathrm{dom~{}}\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b})$, $\tau^{(0)}=0$, and starting from $k=1$ let
$\boldsymbol{d}^{(k-1)}$ $=\boldsymbol{d}(\boldsymbol{p}^{(k-1)};t,\boldsymbol{b}),$ (29)
$\Delta^{(k-1)}$ $=\Delta_{*}(\boldsymbol{p}^{(k-1)};t,\boldsymbol{b}),$
$\tau^{(k)}$ $=\tau_{*}(\boldsymbol{p}^{(k-1)};t,\boldsymbol{b}),$
$\boldsymbol{p}^{(k)}$ $=\boldsymbol{p}^{(k-1)}+\begin{cases}\min\left(\Delta^{(k-1)},1/t\right)\boldsymbol{d}^{(k-1)}&\text{if}\ t>0\\ \Delta^{(k-1)}\boldsymbol{d}^{(k-1)}&\text{if $t=0$ and $\Delta^{(k-1)}<+\infty$},\\ \boldsymbol{0}&\text{otherwise.}\end{cases}$
The slow solution is given piecewise on the intervals $[\tau^{(k-1)},\tau^{(k)})$:
$\boldsymbol{p}(\tau)=\boldsymbol{p}^{(k-1)}+f(\tau-\tau^{(k-1)},t)\boldsymbol{d}^{(k-1)}\ \text{over}\ \tau\in[\tau^{(k-1)},\tau^{(k)}).$ (30)
This gives the explicit solution of the slow system over $[0,\tau^{(k)})$ up to any $k\in\mathbb{N}$. Note that if $t>0$, then the minimum in (29) is attained at $1/t\iff\tau^{(K)}=+\infty$ for some $K\in\mathbb{N}$.
What can be said about the asymptotic limit $k\to+\infty$? The following Theorem asserts the limit converges in finite time, in the sense that there exists some nonnegative integer $K$ such that $\boldsymbol{p}(\tau)=\boldsymbol{p}^{(K)}$ over the interval $[\tau^{(K)},+\infty)$.
Theorem 4.2.
Let $t\geqslant 0$ and $\boldsymbol{p}^{(0)}\in\mathrm{dom~{}}\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b})$. Consider the slow system (12) with initial condition $\boldsymbol{p}(0)=\boldsymbol{p}^{(0)}$, whose solution is given by (30). Then there exists a nonnegative integer $K$ such that $\boldsymbol{d}^{(K)}=\boldsymbol{0}$ and $\boldsymbol{p}(\tau)=\boldsymbol{p}^{(K)}$ on $\tau\in[\tau^{(K)},+\infty)$.
Proof.
We will show there exists a nonnegative integer $K$ for which $\boldsymbol{d}^{(K)}=\boldsymbol{0}$. This suffices because the projection onto the closed, convex set $\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b})$ is unique (see A.17), and so the subgradients are different if and only if their norms are different.
Let $k$ be a positive integer and suppose $\tau^{(j)}<+\infty$ for every $j\in\{1,\dots,k\}$ in (29). On the one hand, Lemma 4.1 implies $\left\|{\boldsymbol{d}^{(j)}}\right\|_{2}<\left\|{\boldsymbol{d}^{(j-1)}}\right\|_{2}$. Hence the descent directions $\{\boldsymbol{d}^{(1)},\dots,\boldsymbol{d}^{(k)}\}$ are obtained from projections landing all on different faces of the convex cone $\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b})$. On the other hand, since $\partial_{\boldsymbol{p}}V(\cdot;t,\boldsymbol{b})$ has at most $4^{n}$ faces and $\boldsymbol{b}\in\mathrm{Im}({\bm{A}})$, there is at least one face on which the projection is zero. Let $K$ denote the smallest such integer in the construction (30). Then $\boldsymbol{d}^{(K)}=\boldsymbol{0}$ and $\boldsymbol{p}(\tau)=\boldsymbol{p}^{(K)}$ on $[\tau^{(K)},+\infty)$. ∎
Remark 4.1 (Drift to infinity, again).
Without the full row rank assumption $\mathrm{rank}({\bm{A}})=m$ (more generally, $\boldsymbol{b}\in\text{Im}(\bm{A})$), there may be some $K\in\mathbb{N}$ for which $\boldsymbol{d}^{(K-1)}\neq 0$ with $\bm{A}^{\top}\boldsymbol{d}^{(K-1)}=\boldsymbol{0}$, so that $\Delta^{(K-1)}=+\infty$ and $\tau^{(K)}=+\infty$. This does not cause an issue when $t>0$ since $\boldsymbol{p}^{(K)}$ remains finite. However, when $t=0$, we have $\lim_{\tau\to+\infty}\left\|{\boldsymbol{p}(\tau)}\right\|_{2}=+\infty$, which means the trajectory of the slow system diverges and there are no feasible solutions to the corresponding (BP) and (dBP) problems.
4.4 Recovery of solutions to basis pursuit denoising
The analysis in Sections 4.14.3 yields the global solution to the slow system (12). Crucially, its asymptotic limit can be computed explicitly from (30) because it converges in finite time, meaning $\lim_{\tau\to+\infty}\boldsymbol{p}(\tau)=\boldsymbol{p}^{(K)}$ for some nonnegative integer $K$. This recovers a pair of primal and dual solutions $(\boldsymbol{x}^{s}(t_{0},\boldsymbol{b}),\boldsymbol{p}^{s}(t,\boldsymbol{b}))$ to BPDN. Indeed, the minimal selection princple implies $\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})=\boldsymbol{p}^{(K)}$, while a primal solution follows from the optimality conditions (6) and Lemma 3.1: Letting $\hat{\boldsymbol{u}}^{(K)}$ denote an optimal solution to the NNLS problem (16), then
$-t\boldsymbol{p}^{(K)}=\boldsymbol{b}-\bm{A}\bm{D}(\boldsymbol{p}^{(K)})\hat{\boldsymbol{u}}^{(K)}\implies\boldsymbol{x}^{s}(t,\boldsymbol{b})=\bm{D}(\boldsymbol{p}^{(K)})\hat{\boldsymbol{u}}^{(K)}.$
4.5 An active set algorithm for basis pursuit denoising
Taken together, our results yield Algorithm 1, presented in the introduction of this work. It computes the trajectory of slow system (12) until the asymptotic limit is reached in finite time. Thanks to (29) and (30), the trajectory is integrated exactly because the descent direction and timestep can be computed using, e.g., active set NNLS or cone projection algorithms as per Remark 3.2 in Section 3.2.
Algorithm 1 requires some input $\boldsymbol{p}^{(0)}\in\mathbb{R}^{m}$ with $\left\|{\bm{A}^{\top}\boldsymbol{p}^{(0)}}\right\|_{\infty}=1$ to serve as the initial condition of the slow system (12). For this, one can take $\boldsymbol{p}^{(0)}=-\boldsymbol{b}/\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}$. In addition, Algorithm 1 is naturally suited for computing regularization paths. To do so, one selects a decreasing sequence of nonnegative hyperparameters $\{t^{(l)}\}_{l=0}^{L}$ and sequentially computes the dual solutions $\{\boldsymbol{p}^{s}(t^{(l)},\boldsymbol{b})\}_{l=0}^{L}$, using $\boldsymbol{p}^{s}(t^{(l)},\boldsymbol{b})$ as the input in Algorithm 1 for computing $\boldsymbol{p}^{s}(t^{(l+1)},\boldsymbol{b})$. Again, one can start with $t^{(0)}=\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}$ and $\boldsymbol{p}^{s}(t^{(0)},\boldsymbol{b})=-\boldsymbol{b}/t^{(0)}$.
5 Perturbation analysis of the slow system
Sections 34 describes how BPDN is characterized from the dynamics of the slow system (12) and its slow solution (30). In particular, the analysis yields an exact algorithm that computes a pair of primal and dual solutions to BPDN for fixed hyperparameter $t$ and data $\boldsymbol{b}$. What happens, however, to the dynamics of the slow system (12) when the hyperparameter or data or both are perturbed? This matter in practice if, for example, one seeks to assess robustness of solutions with respect to the hyperparameter and data [21, 46, 28]. Another example is if the desired hyperparameter is unknown. In this case, one typically compute a regularization path or the entire solution path $t\mapsto(\boldsymbol{x}^{s}(t,\boldsymbol{b}),\boldsymbol{p}^{s}(t,\boldsymbol{b}))$ and then select the desired hyperparameter, e.g., based on the sparsity of the solution or data-driven methods [30, 50, 25, 22, 11]. In any case, the perturbation analysis of BPDN is crucial to tackle these questions. In the context of minimal selection principle, this amounts to analyzing the dependence on parameters of the slow system and its slow solution.
Motivated by these matters, we characterize here how the dynamics of the slow system and its slow solution behave under perturbations on the hyperparameter $t$ and data $\boldsymbol{b}$. First, we present in Section 5.1 how the descent direction of the slow system behaves under perturbations. We then use this in conjunction with the minimal selection principle in Section 5.2 to describe the local dependence on the parameters $t$ and $\boldsymbol{b}$ of the slow solution and, therefore, the primal and dual solutions to BPDN. As we argue in Section 5.3, this local dependence leads to a possibly non-unique local continuation of the primal solution to BPDN. This non-uniqueness issue, a well-known problem in the literature [63], arises from the non-uniqueness of solutions to an NNLS problem, as also previously reported by [11]. Finally, we present in Section 5.4 the global dependence on parameters of the slow solution, naturally yielding a rigorous homotopy algorithm, similar to the one proposed in [11] but based firmly on the minimal selection principle, for computing the solution path to BPDN.
5.1 Perturbations of the descent direction
We first describe how the descent direction $\boldsymbol{d}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})$ changes under perturbations in the hyperparameter and data. This will be used in Section 5.2 to characterize the local dependence on parameters of the slow solution (30).
Lemma 5.1.
Let $t_{0}\geqslant 0$, let $\boldsymbol{p}_{0}\in\mathrm{dom~{}}V(\cdot;t_{0},\boldsymbol{b})$, and let $\hat{\boldsymbol{u}}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})$ be a global minimum of the NNLS problem (16) so that $\boldsymbol{d}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})=\bm{A}\bm{D}(\boldsymbol{p}_{0})\hat{\boldsymbol{u}}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})-(\boldsymbol{b}+t_{0}\boldsymbol{p}_{0})$. In addition, let $\delta_{0}\in\mathbb{R}$ be such that $t_{0}+\delta_{0}\geqslant 0$ and let $\boldsymbol{\delta}\boldsymbol{b}\in\mathbb{R}^{m}$.
  • (i)
    The perturbed descent direction $\boldsymbol{d}(\boldsymbol{p}_{0};t_{0}+\delta_{0},\boldsymbol{b}+\boldsymbol{\delta}\boldsymbol{b})$ is given by
    $\boldsymbol{d}(\boldsymbol{p}_{0};t_{0}+\delta_{0},\boldsymbol{b}+\boldsymbol{\delta}\boldsymbol{b})=\boldsymbol{d}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})+\bm{A}\bm{D}(\boldsymbol{p}_{0})\hat{\boldsymbol{v}}(\boldsymbol{p}_{0};t,\boldsymbol{b},\delta_{0},\boldsymbol{\delta}\boldsymbol{b})-\boldsymbol{\delta}\boldsymbol{b}-\delta_{0}\boldsymbol{p}_{0},$ (31)
    where
    $\hat{\boldsymbol{v}}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b},\delta_{0},\boldsymbol{\delta}\boldsymbol{b})\in\operatorname*{arg\,min}_{\boldsymbol{v}\in\mathbb{R}^{n}}\left\|{\boldsymbol{d}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})+\bm{A}\bm{D}(\boldsymbol{p}_{0})\boldsymbol{v}-\boldsymbol{\delta}\boldsymbol{b}-\delta_{0}\boldsymbol{p}_{0}}\right\|_{2}^{2}$ (32)
    subject $\qquad\qquad\text{subject to}\ \begin{cases}\boldsymbol{v}_{\mathcal{E}(\boldsymbol{p}_{0})}&\geqslant-\hat{\boldsymbol{u}}_{\mathcal{E}(\boldsymbol{p}_{0})}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b}),\\ \boldsymbol{v}_{\mathcal{E}^{\mathsf{C}}(\boldsymbol{p}_{0})}&=\boldsymbol{0}.\end{cases}$
  • (ii)
    (Linear perturbations in hyperparameter) Suppose $\boldsymbol{d}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})=\boldsymbol{0}$, let $\delta_{0}=t-t_{0}$ with $t\in[0,t_{0}]$, and let $\delta_{0}\boldsymbol{b}=-(t_{0}-t)\boldsymbol{q}$ with $\boldsymbol{q}\in\mathbb{R}^{m}$. Then (31) and (32) refine to
    $\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q})=(1-t/t_{0})\left(\bm{A}\bm{D}(\boldsymbol{p}_{0})\hat{\boldsymbol{v}}(\boldsymbol{p}_{0};t_{0},t,\boldsymbol{b},\boldsymbol{q})+t_{0}(\boldsymbol{p}_{0}+\boldsymbol{q})\right)$ (33)
    where
    $\hat{\boldsymbol{v}}(\boldsymbol{p}_{0};t_{0},t,\boldsymbol{b},\boldsymbol{q})\in\operatorname*{arg\,min}_{\boldsymbol{v}\in\mathbb{R}^{n}}\left\|{\bm{A}\bm{D}(\boldsymbol{p}_{0})\boldsymbol{v}+t_{0}(\boldsymbol{p}_{0}+\boldsymbol{q})}\right\|_{2}^{2}$ (34)
    subject $\text{subject to}\ \begin{cases}&\boldsymbol{v}_{j}\geqslant-\hat{\boldsymbol{u}}_{j}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})/(1-t/t_{0})\ \text{if $j\in\mathcal{E}(\boldsymbol{p}_{0})$},\\ &\boldsymbol{v}_{\mathcal{E}^{\mathsf{C}}(\boldsymbol{p}_{0})}=\boldsymbol{0}.\end{cases}$
Proof.
See Appendix B.3 for the proof. ∎
5.2 Local dependence on parameters of the slow solution
In the stationary case $\boldsymbol{d}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})=\boldsymbol{0}$ with $t_{0}>0$, Lemma 5.1 characterizes implicitly how the primal and dual solutions to BPDN change under perturbations in the hyperparameter and data. Here, we show how these perturbations characterize these changes explicitly.
Proposition 5.1.
Let $t_{0}>0$, $\boldsymbol{q}\in\mathbb{R}^{m}$, and $(\boldsymbol{x}^{s}(t_{0},\boldsymbol{b}),\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$ denote a pair of primal and dual solutions to BPDN at hyperparameter $t_{0}$ and data $\boldsymbol{b}$. In addition, let
$\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})\in\operatorname*{arg\,min}_{\boldsymbol{v}\in\mathbb{R}^{n}}\left\|{\bm{A}\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\boldsymbol{v}+t_{0}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})+\boldsymbol{q})}\right\|_{2}^{2}$ (35)
subject $\text{subject to}\ \begin{cases}\boldsymbol{v}_{j}&\geqslant 0\ \text{if $j\in\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$ and $\boldsymbol{x}^{s}_{j}(t_{0},\boldsymbol{b})=0$},\\ \boldsymbol{v}_{j}&=0\ \text{if $j\in\mathcal{E}^{\mathsf{C}}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$},\end{cases}$
and define
$\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})\coloneqq\bm{A}\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})+t_{0}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})+\boldsymbol{q}),$ (36)
$C^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})\coloneqq$
$\quad\inf_{j\in\{1,\dots,n\}}\left\{\frac{\mathrm{sgn}{\left(\langle\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\bm{A}^{\top}\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q}),\boldsymbol{e}_{j}\rangle\right)}-\langle\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\bm{A}^{\top}\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}),\boldsymbol{e}_{j}\rangle)}{\langle\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\bm{A}^{\top}\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q}),\boldsymbol{e}_{j}\rangle}\right\},$
$T_{+}(t_{0},\boldsymbol{b},\boldsymbol{q})\coloneqq\frac{t_{0}}{1+t_{0}C^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})},$
$T_{-}(t_{0},\boldsymbol{b},\boldsymbol{q},\hat{\boldsymbol{v}}^{s})\coloneqq t_{0}\left(1-\inf_{\begin{subarray}{c}j\in\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\\ \boldsymbol{x}^{s}_{j}(t_{0},\boldsymbol{b})\neq 0\\ \hat{\boldsymbol{v}}_{j}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})\leqslant-|\boldsymbol{x}_{j}^{s}(t_{0},\boldsymbol{b})|\end{subarray}}\frac{|\boldsymbol{x}_{j}^{s}(t_{0},\boldsymbol{b})|}{|\hat{\boldsymbol{v}}^{s}_{j}(t_{0},\boldsymbol{b},\boldsymbol{q})|}\right),$
$t_{1}\coloneqq\max\left(T_{-}(t_{0},\boldsymbol{b},\boldsymbol{q},\hat{\boldsymbol{v}}^{s}),T_{+}(t_{0},\boldsymbol{b},\boldsymbol{q})\right).$
Then, we have the following:
  • (i)
    There holds $0\leqslant t_{1}<t_{0}$, and the primal and dual solutions to basis pursuit denoising at hyperparameter $t\in[t_{1},t_{0}]$ and data $\boldsymbol{b}-(t_{0}-t)\boldsymbol{q}$ are given by
    $\boldsymbol{x}^{s}(t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q})$ $=\boldsymbol{x}^{s}(t_{0},\boldsymbol{b})+\left(1-\frac{t}{t_{0}}\right)\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q}),$ (37)
    $\boldsymbol{p}^{s}(t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q})$ =$=\begin{cases}&\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})+\left(\frac{1}{t}-\frac{1}{t_{0}}\right)\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})\quad\text{if $t_{1}(t_{0},\boldsymbol{b},\boldsymbol{q})>0$},\\ &\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})\quad\text{otherwise.}\end{cases}$
  • (ii)
    (Monotonicity) Assume $\boldsymbol{q}=-\bm{A}\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})\boldsymbol{w}/t_{0}$ for some nonnegative $\boldsymbol{w}\in\mathbb{R}^{n}$. Then
    $\begin{gathered}V(\boldsymbol{p}^{s}(t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q});t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q})\leqslant V(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\\ \left\|{\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})}\right\|_{2}\leqslant\left\|{\boldsymbol{p}^{s}(t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q})}\right\|_{2}.\end{gathered}$ (38)
    If also $\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})\neq\boldsymbol{0}$, then both inequalities in (38) are strict.
Proof.
See Appendix B.4 for the proof. ∎
Remark 5.1.
The numbers $T_{+}(t_{0},\boldsymbol{b},\boldsymbol{q})$ and $T_{-}(t_{0},\boldsymbol{b},\boldsymbol{q},\hat{\boldsymbol{v}}^{s})$ identify potential “kinks” in the piecewise linear dependence of the primal and dual solutions on $t$. If $t_{1}=T_{+}(t_{0},\boldsymbol{b},\boldsymbol{q})>0$, then there is at least one index $j\in\{\{1,\dots,n\}:\boldsymbol{x}^{s}_{j}(t_{0},\boldsymbol{b})=0\}$ that joins the set $\mathcal{E}(\boldsymbol{p}^{s}(t_{1},\boldsymbol{b}-(t_{0}-t_{1})\boldsymbol{q}))$, and this index is new if $j\in\mathcal{E}^{\mathsf{C}}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$. If $t_{1}=T_{-}(t_{0},\boldsymbol{b},\boldsymbol{q},\hat{\boldsymbol{v}}^{s})>0$, then there is at least one index $j\in\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$ such that $\boldsymbol{x}^{s}_{j}(t_{1},\boldsymbol{b}-(t_{0}-t_{1})\boldsymbol{q})=\boldsymbol{0}$. This index may leave the equicorrelation set $\mathcal{E}(\boldsymbol{p}^{s}(t_{1},\boldsymbol{b}-(t_{0}-t_{1})))$ beyond $t>t_{1}$.
5.3 Non-uniqueness in the local dependence on parameters
The local continuation of $(\boldsymbol{x}^{s}(t_{0},\boldsymbol{b}),\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$ depends on the solution $\boldsymbol{v}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})$ and residual vector $\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})$ of the NNLS problem (35). Since the residual vector is always unique, both the number $T_{+}(t_{0},\boldsymbol{b},\boldsymbol{q})$ and the continuation of $\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})$ are also unique. The NNLS solution, however, is generally not unique. This leads to complications because both the number $T_{-}(t_{0},\boldsymbol{b},\boldsymbol{q},\hat{\boldsymbol{v}}^{s})$ and continuation of $\boldsymbol{x}^{s}(t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q})$ beyond $t<t_{0}$ may not be unique [^2] This problem was pointed out earlier by Bringmann et al. [11] in a different context, where the authors constructed a homotopy algorithm directly from the KKT conditions of BPDN. Their NNLS problem (Equation 3.2. on page 6) is equivalent to (35) with $\boldsymbol{q}=\boldsymbol{0}$.. This is the well-known problem of non-uniqueness of primal solutions to BPDN [63].
To understand how the non-uniqueness of the NNLS problem (35) arises, it helps to identify when the solution is unique. If the least-squares solution
$\hat{\boldsymbol{v}}^{\text{LSQ}}=t_{0}(\bm{A}^{\top}_{\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})}\bm{A}_{\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))})^{-1}\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\bm{A}^{\top}_{\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})+\boldsymbol{q})$
satisfies $\hat{\boldsymbol{v}}^{\text{LSQ}}_{j}\geqslant 0$ for every $j\in\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$ with $\boldsymbol{x}_{j}^{s}(t_{0},\boldsymbol{b})=0$ and $\bm{A}_{\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}$ has full column rank, then it is the unique solution. If instead $\bm{A}_{\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}$ has full row rank, then $\hat{\boldsymbol{v}}^{\text{LSQ}}$ is the unique solution with least $\ell_{2}$ norm. Thus one possibility is to select the solution with least $\ell_{2}$-norm. However, in practice we have found that it was unnecessary; see Remark 5.2.
5.4 Global dependence on parameters of the slow solution
From a dynamical systems perspective, Lemma 5.1 and Proposition 5.1 describe the local continuation of the asymptotic limit of the slow system (12) in terms of the hyperparameter and data. This continuation is local because it holds on an interval $t\in[t_{1},t_{0}]$. As discussed in Section 5.3, the continuation is not unique if the solution to the NNLS problem (35) is not unique. Nevertheless, if $t_{1}\neq 0$, we can apply Proposition 5.1 again to extend the local continuation to an interval $[t_{2},t_{0}]$ with $0\leqslant t_{2}<t_{1}$.
We now apply Proposition 5.1 repeatedly with $\boldsymbol{q}=\boldsymbol{0}$ throughout, starting from $t^{(0)}=\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}$, $\boldsymbol{x}^{(0)}=\boldsymbol{0}$, and $\boldsymbol{p}^{(0)}=-\boldsymbol{b}/\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}$. Doing so $k$ times yields a global continuation of the asymptotic limit of the slow system (12), hence a global continuation of the pair of primal and dual solutions $(\boldsymbol{x}^{s}(t,\boldsymbol{b}),\boldsymbol{p}^{s}(t,\boldsymbol{b}))$ on $t\in[t^{(k)},t^{(0)}]$.
As discussed in Section 5.3, the continuation in the primal solution $\boldsymbol{x}^{s}(t,\boldsymbol{b})$ is not unique. Choosing at each $k$ the NNLS solution to problem (35) with least $\ell_{2}$ norm, we obtain Algorithm 2 below. It is nearly identical to the generalized homotopy algorithm proposed in Bringmann et al. [11], with the exception of the breakpoints, which here are simpler and more explicit. Algorithm 2 yields a solution path of BPDN and converges in finite time.
Input : A matrix $\bm{A}\in\mathbb{R}^{m\times n}$ and a vector $\boldsymbol{b}\in\mathrm{Im}({\bm{A}})\setminus{\{\boldsymbol{0}\}}$.
Output : A finite sequence $\left\{t^{(k)},\boldsymbol{x}^{(k)},\boldsymbol{p}^{(k)}\right\}_{k=0}^{K}$ specifying a solution path to the basis pursuit denoising problem.
1 Set $t^{(0)}=\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}$, $\boldsymbol{x}^{(0)}=\boldsymbol{0}$ and $\boldsymbol{p}^{(0)}=-\boldsymbol{b}/\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}$;
2 for $k=1$ until convergence do
3       Compute $\mathcal{E}^{(k-1)}=\{j\in\{1,\dots,n\}:|\langle-\bm{A}^{\top}\boldsymbol{p},\boldsymbol{e}_{j}\rangle|=1\}$;
4       Compute $\bm{D}^{(k-1)}=\mathrm{diag}\left(\mathrm{sgn}{(-\bm{A}^{\top}\boldsymbol{p}^{(k-1)}})\right)$;
5       Compute the solution $\hat{\boldsymbol{v}}^{(k-1)}\in\operatorname*{arg\,min}_{\boldsymbol{v}\in\mathbb{R}^{n}}\left\|{\bm{A}_{\mathcal{E}^{(k-1)}}\bm{D}_{\mathcal{E}^{(k-1)}}^{(k-1)}\boldsymbol{v}+t^{(k-1)}\boldsymbol{p}^{(k-1)}}\right\|_{2}^{2}$ subject to
$\begin{cases}\boldsymbol{v}_{j}\geqslant 0,\,&\text{if $j\in\mathcal{E}^{(k-1)}$ and $\boldsymbol{x}_{j}^{(k-1)}=0$}\\ \boldsymbol{v}_{j}=0,\,&\text{if $j\in\left(\mathcal{E}^{(k-1)}\right)^{\mathsf{C}}$}\end{cases}$
with minimal $\ell_{2}$ norm;
6       Compute $\boldsymbol{\xi}^{(k-1)}=\bm{A}\bm{D}(\boldsymbol{p}^{(k-1)})\hat{\boldsymbol{v}}^{(k-1)}+t^{(k-1)}\boldsymbol{p}^{(k-1)}$ ;
7       Compute the numbers $T_{-}(t^{(k-1)},\boldsymbol{b},\boldsymbol{0},\hat{\boldsymbol{v}}^{(k-1)})$ and $T_{+}(t^{(k-1)},\boldsymbol{b},\boldsymbol{0})$ from (36);
8       Update $t^{(k)}=\max(T_{-}(t^{(k-1)},\boldsymbol{b},\boldsymbol{0},\hat{\boldsymbol{v}}^{(k-1)}),T_{+}(t^{(k-1)},\boldsymbol{b},\boldsymbol{0}))$;
9       Update $\boldsymbol{x}^{(k)}=\boldsymbol{x}^{(k-1)}+\left(1-t^{(k)}/t^{(k-1)}\right)\bm{D}(\boldsymbol{p}^{(k-1)})\hat{\boldsymbol{v}}^{(k-1)}$;
10       if $t^{(k)}=0$ then
11             $\boldsymbol{p}^{(k)}=\boldsymbol{p}^{(k-1)}$ ;
12             break // The algorithm has converged ;
13
14       else
15             Update $\boldsymbol{p}^{(k)}=\boldsymbol{p}^{(k-1)}+(1/t^{(k)}-1/t^{(k-1)})\boldsymbol{\xi}^{(k-1)}$;
16
17
18 end for
Algorithm 2 Homotopy algorithm for computing the primal/dual solution paths to the basis pursuit denoising problem.
Theorem 5.1.
The global continuation method described in Algorithm 2 is correct and converges in finite time, that is, there exists a nonnegative integer $K$ such that $t^{(K)}=0$.
Proof.
Correctness follows immediately from Proposition 5.1 and the discussion in Section 5.3. The proof of convergence in finite time is identical to that of Theorem 4.2 in Bringmann et al. [11] and is omitted. ∎
Remark 5.2.
In practice, we have found that Algorithm 2 always converged when using the Lawson–Hanson algorithm [41] or Meyers’s algorithm [47] to compute a solution to the NNLS problem on line 5 without explicitly finding the minimal $\ell_{2}$-norm solution. In particular, this worked on Mairal’s worst-case exponential complexity example [45] or the pathological example in [11, Proposition 4.1].
We omitted in Theorem 5.1 the global dependence on the data $\boldsymbol{b}$ via linear perturbations $\boldsymbol{b}-(t_{0}-t)\boldsymbol{q}$ with $t\leqslant t_{0}$ and $\boldsymbol{q}\in\mathbb{R}^{m}$. In the next section, we consider these perturbations and propose a greedy approach that continuously change hyperparameter and data to compute a pair of feasible primal and dual solutions to BP in strongly polynomial time.
6 Greedy approach for feasibility of basis pursuit
Sections 5.15.3 describes the local dependence of the asymptotic limit of the slow system (12) on the hyperparameter $t$ and data $\boldsymbol{b}$. The global dependence on the data was omitted in Section 5.4 to provide a characterization of the global dependence of the asymptotic limit of the slow system on the hyperparameter $t$ only, thereby yielding Algorithm 2, a generalized homotopy algorithm, nearly identical to the one proposed in [11] but based firmly on the minimal selection principle. In this section, we analyze the global dependence of the asymptotic limit of the slow system (12) on both hyperparameter and data.
This may seem a strange thing to do. After all, in BPDN the data is typically fixed. Why do this in the first place? Well, because it can be useful: perturbations in the data can be added to compute a feasible solution to (BP) in strongly polynomial time. This provides either the optimal solution if BP has a unique solution or it provides an approximation that can be used as a warm start for Algorithm 1 with $t=0$ to compute an optimal BP solution.
The idea, detailed in Sections 6.16.3, is to introduce a perturbation $\boldsymbol{q}$ in the NNLS problem (35) to ensure the solution $\hat{\boldsymbol{v}}^{s}$ is the (unique) least-squares solution and the breakpoint $T_{-}(t_{0},\boldsymbol{b},\boldsymbol{q},\hat{\boldsymbol{v}}^{s})\leqslant 0$, that is, no element of the equicorrelation set is removed. Doing this repeatedly, the equicorrelation set grows by at least one index while simultaneously marking linearly independent columns of the matrix $\bm{A}$. The latter holds because the dependence on parameters of the slow solutions admits, under certain condition, a greedy property.
6.1 Existence of a greedy property
Our starting point is the monotonic property stated in Proposition 5.1(ii), which holds for data perturbations $\boldsymbol{q}=-\bm{A}\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\boldsymbol{w}/t_{0}$ with nonnegative $\boldsymbol{w}\in\mathbb{R}^{n}$. We show that certain choices of $\boldsymbol{w}$ ensure the local dependence on parameters satisfies a greedy property.
Proposition 6.1.
Let $(\boldsymbol{x}^{s}(t_{0},\boldsymbol{b}),\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$ denote a pair of primal and dual solutions to BPDN at hyperparameter $t_{0}>0$ and data $\boldsymbol{b}\in\mathbb{R}^{m}$. Let $\hat{\boldsymbol{v}}^{\text{LSQ}}(t_{0},\boldsymbol{b})\in\mathbb{R}^{n}$ denote the vector
$\begin{cases}\hat{\boldsymbol{v}}^{\text{LSQ}}_{\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}(t_{0},\boldsymbol{b})&=t_{0}\left(\bm{A}^{\top}_{\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}\bm{A}_{\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}\right)^{-1}\boldsymbol{1}_{\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))},\\ \hat{\boldsymbol{v}}^{\text{LSQ}}_{\mathcal{M}^{\mathsf{C}}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}(t_{0},\boldsymbol{b})&=\boldsymbol{0},\end{cases}$ (39)
where
$\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\in\sup_{|\mathcal{M}|}\left\{\text{$\mathcal{M}\subset\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$ : s.t. $\left\{\bm{A}\boldsymbol{e}_{j}\right\}_{j\in\mathcal{M}}$ are linearly independent}\right\}.$ (40)
In addition, consider the NNLS problem (35) with the greedy selection
$\boldsymbol{q}=-\bm{A}\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\boldsymbol{w}/t_{0},$ (41)
where $\boldsymbol{w}\in\mathbb{R}^{n}$ satisfies
$\begin{cases}\boldsymbol{w}_{\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}&\geqslant-\bm{D}_{\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}\boldsymbol{x}^{s}_{\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}(t_{0},\boldsymbol{b})-\hat{\boldsymbol{v}}^{\text{LSQ}}(t_{0},\boldsymbol{b}),\\ \boldsymbol{w}_{\mathcal{M}^{\mathsf{C}}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}&=-\bm{D}_{\mathcal{M}^{\mathsf{C}}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}\boldsymbol{x}^{s}_{\mathcal{M}^{\mathsf{C}}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}(t_{0},\boldsymbol{b}).\\ \end{cases}$ (42)
Then, we have the following.
  • (i)
    The set of solutions to problem (35) includes $\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b},\boldsymbol{w})=\hat{\boldsymbol{v}}^{\text{LSQ}}(t_{0},\boldsymbol{b})+\boldsymbol{w}$, and the residual vector $\boldsymbol{\xi}^{s}$ and breakpoint $T_{+}$ in (36) are independent of $\boldsymbol{w}$. Moreover,
    $\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b})=\bm{A}_{\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}\bm{D}_{\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}\hat{\boldsymbol{v}}^{\text{LSQ}}(t_{0},\boldsymbol{b})+t_{0}\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}),$ (43)
    $\bm{A}^{\top}_{\mathcal{E}{(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}}\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b})=\boldsymbol{0}\quad\text{and}\quad T_{-}(t_{0},\boldsymbol{b},\boldsymbol{w},\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b},\boldsymbol{w}))\leqslant 0\leqslant T_{+}(t_{0},\boldsymbol{b}).$
  • (ii)
    For every $t\in(T_{+}(t_{0},\boldsymbol{b}),t_{0}]$, we have the set equality
    $\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))=\mathcal{E}\left(\boldsymbol{p}^{s}(t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q})\right),$ (44)
    and if $T_{+}(t_{0},\boldsymbol{b})>0$, we have the strict inclusion
    $\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\subsetneq\mathcal{E}\left(\boldsymbol{p}^{s}(T_{+}(t_{0},\boldsymbol{b}),\boldsymbol{b}-(t_{0}-T_{+}(t_{0},\boldsymbol{b}))\boldsymbol{q})\right).$ (45)
  • (iii)
    (Greedy property.) Let $j$ be in the symmetric difference of the sets in (45). The column vector $\bm{A}\boldsymbol{e}_{j}$ is linearly independent from the column vectors $\{\bm{A}\boldsymbol{e}_{k}\}_{k\in\mathcal{E}\left(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})\right)}$.
Proof.
See Appendix B.5 for the proof. ∎
6.2 Specifying the greedy selection
The greedy selection (41) works for any $\boldsymbol{w}\in\mathbb{R}^{n}$ satisfying the set of mixed equalities and inequalities in (42). What $\boldsymbol{w}$ should be chosen? One possibility it to select the one with minimal $\ell_{2}$-norm. That is, we choose
$\boldsymbol{w}^{\text{min}}=\operatorname*{arg\,min}_{\boldsymbol{w}\in\mathbb{R}^{n}}\left\|{\boldsymbol{w}}\right\|_{2}^{2}\ \text{subject to}$
$\boldsymbol{w}_{\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}\geqslant-\hat{\boldsymbol{v}}^{\text{LSQ}}_{\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}(t_{0},\boldsymbol{b})-\bm{D}_{\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}\boldsymbol{x}^{s}_{\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}(t_{0},\boldsymbol{b}),$
$\boldsymbol{w}_{\mathcal{M}^{\mathsf{C}}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}=-\bm{D}_{\mathcal{M}^{\mathsf{C}}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}\boldsymbol{x}^{s}_{\mathcal{M}^{\mathsf{C}}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}(t_{0},\boldsymbol{b}).$
The solution to this problem exists, is unique, and has the closed-form expression
$\begin{cases}\boldsymbol{w}^{\text{min}}_{\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}(t_{0},\boldsymbol{b})&=\min(\boldsymbol{0},-\bm{D}_{\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}\boldsymbol{x}^{s}_{\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}(t_{0},\boldsymbol{b})-\hat{\boldsymbol{v}}^{\text{LSQ}}_{\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}(t_{0},\boldsymbol{b})),\\ \boldsymbol{w}^{\text{min}}_{\mathcal{M}^{\mathsf{C}}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})}(t_{0},\boldsymbol{b})&=-\bm{D}_{\mathcal{M}^{\mathsf{C}}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}\boldsymbol{x}^{s}_{\mathcal{M}^{\mathsf{C}}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}(t_{0},\boldsymbol{b}).\end{cases}$ (46)
In this case, the solution $\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b},\boldsymbol{w}^{\text{min}})$ to problem (35) equals
$\begin{cases}\boldsymbol{v}^{s}_{\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}(t_{0},\boldsymbol{b},\boldsymbol{w}^{\text{min}})&=\min(\hat{\boldsymbol{v}}^{\text{LSQ}}_{\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}(t_{0},\boldsymbol{b}),-\bm{D}_{\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}\boldsymbol{x}^{s}_{\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}(t_{0},\boldsymbol{b})),\\ \boldsymbol{v}^{s}_{\mathcal{M}^{\mathsf{C}}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}(t_{0},\boldsymbol{b},\boldsymbol{w}^{\text{min}})&=-\bm{D}_{\mathcal{M}^{\mathsf{C}}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}\boldsymbol{x}^{s}_{\mathcal{M}^{\mathsf{C}}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}(t_{0},\boldsymbol{b}).\end{cases}$ (47)
6.3 A greedy algorithm for computing feasible solutions
We now combine Propositions 5.1 and 6.1 to obtain Algorithm 3. This greedy algorithm computes a pair of feasible primal and dual solutions to (BP) and (dBP) in strongly polynomial time, as Theorem 6.1 below demonstrates.
Input : A matrix $\bm{A}\in\mathbb{R}^{m\times n}$ and a vector $\boldsymbol{b}\in\mathrm{Im}({\bm{A}})\setminus{\{\boldsymbol{0}\}}$.
Output : A pair of vectors $(\boldsymbol{x}_{f},\boldsymbol{p}_{f})\in\mathbb{R}^{n}\times\mathbb{R}^{m}$ where $\boldsymbol{x}_{f}$ has at most $m$ nonzero components, $\bm{A}\boldsymbol{x}_{f}=\boldsymbol{b}$, and $\left\|{-\bm{A}^{\top}\boldsymbol{p}_{f}}\right\|_{\infty}\leqslant 1$.
1 Set $\boldsymbol{b}^{(0)}=\boldsymbol{b}$, $t^{(0)}=\left\|{\bm{A}^{\top}\boldsymbol{b}^{(0)}}\right\|_{\infty}$, $\boldsymbol{x}^{(0)}=\boldsymbol{0}$, $\boldsymbol{p}^{(0)}=-\boldsymbol{b}^{(0)}/t^{(0)}$, and $\mathcal{M}^{(-1)}=\{1,\dots,n\}$;
2 for $k=1$ until convergence do
3       Compute $\mathcal{E}^{(k-1)}=\{j\in\{1,\dots,n\}:|\langle-\bm{A}^{\top}\boldsymbol{p},\boldsymbol{e}_{j}\rangle|=1\}$;
4       Compute $\mathcal{M}^{(k-1)}\in\sup_{|\mathcal{M}|}\left\{\text{$\mathcal{M}\subset\mathcal{E}(\boldsymbol{p}^{(k-1)})\cap\mathcal{M}^{(k-2)}$ : s.t. $\{\bm{A}\boldsymbol{e}_{j}\}_{j\in\mathcal{M}}$ are linearly independent}\right\}$;
5       Compute $\bm{D}^{(k-1)}=\mathrm{diag}\left(\mathrm{sgn}{(-\bm{A}^{\top}\boldsymbol{p}^{(k-1)}})\right)$;
6       Compute $\hat{\boldsymbol{v}}^{\text{LSQ}}(t^{(k-1)},\boldsymbol{b}^{(k-1)})$ using Equation (39);
7       Compute $\boldsymbol{w}^{(k-1)}=\boldsymbol{w}^{\text{min}}(t^{(k-1)},\boldsymbol{b}^{(k-1)})$ using Equation (46);
8       Compute $\boldsymbol{q}^{(k-1)}=-\bm{A}\bm{D}^{(k-1)}\boldsymbol{w}^{(k-1)}/t^{(k-1)}$;
9       Compute $\hat{\boldsymbol{v}}^{(k-1)}=\hat{\boldsymbol{v}}^{\text{LSQ}}(t^{(k-1)},\boldsymbol{b}^{(k-1)})+\boldsymbol{w}^{(k-1)}$;
10       Compute $\boldsymbol{\xi}^{(k-1)}=t^{(k-1)}\left(\bm{A}_{\mathcal{M}(\boldsymbol{p}^{(k-1)})}\bm{D}^{(k-1)}_{\mathcal{M}(\boldsymbol{p}^{(k-1)})}\hat{\boldsymbol{v}}^{\text{LSQ}}(t^{(k-1)},\boldsymbol{b}^{(k-1)})+\boldsymbol{p}^{(k-1)}\right)$;
11       Compute $C^{(k-1)}=\min_{j\in\mathcal{E}^{\mathsf{C}}{(\boldsymbol{p}^{(k-1)})}}\left\{\frac{\mathrm{sgn}\left(\langle\bm{D}^{(k-1)}\bm{A}^{\top}\boldsymbol{\xi}^{(k-1)},\boldsymbol{e}_{j}\rangle\right)-\langle\bm{D}^{(k-1)}\bm{A}^{\top}\boldsymbol{p}^{(k-1)},\boldsymbol{e}_{j}\rangle)}{\langle\bm{D}^{(k-1)}\bm{A}^{\top}\boldsymbol{\xi}^{(k-1)},\boldsymbol{e}_{j}\rangle}\right\}$;
12       Update $t^{(k)}=t^{(k-1)}/(1+t^{(k-1)}C^{(k-1)})$;
13       Update $\boldsymbol{b}^{(k)}=\boldsymbol{b}^{(k-1)}+\left(1-\frac{t^{(k)}}{t^{(k-1)}}\right)\bm{A}_{\mathcal{E}(\boldsymbol{p}^{(k-1)})}\bm{D}^{(k-1)}_{\mathcal{E}(\boldsymbol{p}^{(k-1)})}\boldsymbol{w}_{\mathcal{E}(\boldsymbol{p}^{(k-1)})}^{(k-1)}$;
14       Update $\boldsymbol{x}^{(k)}=\boldsymbol{x}^{(k-1)}+\left(1-\frac{t^{(k)}}{t^{(k-1)}}\right)\bm{D}^{(k-1)}\boldsymbol{v}^{(k-1)}$;
15       if $t^{(k)}=0$ then
16             Compute $\boldsymbol{x}_{f}=\boldsymbol{x}^{(k)}-\sum_{l=0}^{k-1}\left(1-t^{(k)}/t^{(k-1)}\right)\bm{D}^{(k-1)}\boldsymbol{w}^{(k-1)}$;
17             Compute $\boldsymbol{p}_{f}=\boldsymbol{p}^{(k-1)}$ ;
18             break // The algorithm has converged ;
19
20       else
21             Update $\boldsymbol{p}^{(k)}=\boldsymbol{p}^{(k-1)}+(1/t^{(k)}-1/t^{(k-1)})\boldsymbol{\xi}^{(k-1)}$;
22
23
24 end for
Algorithm 3 Greedy algorithm for computing a pair of feasible primal/dual solutions to the basis pursuit problem.
Theorem 6.1.
After initialization, Algorithm 3 converges in at most $m$ iterations. That is, there exists an integer $K\leqslant m$ such that $t^{(K)}=0$, $\bm{A}\boldsymbol{x}^{(K)}=\boldsymbol{b}^{(K)}$ and
$\boldsymbol{x}^{(K)}$ $\in\operatorname*{arg\,min}_{\boldsymbol{x}\in\mathbb{R}^{n}}\,\left\|{\boldsymbol{x}}\right\|_{1}\quad\text{subject to}\quad\bm{A}\boldsymbol{x}=\boldsymbol{b}^{(K)},$ (48)
$\boldsymbol{p}^{(K)}$ $\in\operatorname*{arg\,min}_{\boldsymbol{p}\in\mathbb{R}^{m}}\,\langle\boldsymbol{p},\boldsymbol{b}^{(K)}\rangle\quad\text{subject to}\quad\left\|{-\bm{A}^{\top}\boldsymbol{p}}\right\|_{\infty}\leqslant 1,$
where
$\boldsymbol{b}^{(K)}=\boldsymbol{b}+\sum_{k=1}^{K}\left(1-t^{(k)}/t^{(k-1)}\right)\bm{A}_{\mathcal{E}(\boldsymbol{p}^{(k-1)})}\bm{D}^{(k-1)}_{\mathcal{E}(\boldsymbol{p}^{(k-1)})}\boldsymbol{w}^{(k-1)}_{\mathcal{E}(\boldsymbol{p}^{(k-1)})}.$ (49)
Moreover, the vectors $(\boldsymbol{x}_{f},\boldsymbol{p}_{f})\in\mathbb{R}^{n}\times\mathbb{R}^{m}$ computed on lines 16-17, namely
$\boldsymbol{x}_{f}$ $=\boldsymbol{x}^{(K)}-\sum_{k=1}^{K}\left(1-t^{(k)}/t^{(k-1)}\right)\bm{D}^{(k-1)}\boldsymbol{w}^{(k-1)}\ \text{and}\ \boldsymbol{p}_{f}=\boldsymbol{p}^{(K-1)},$ (50)
are a pair of feasible primal and dual solutions to (BP) and (dBP).
Proof.
Line 1 initializes the primal and dual solutions at hyperparameter $t^{(0)}=||\bm{A}^{\top}\boldsymbol{b}^{(0)}||_{\infty}$, the smallest hyperparameter for which $(\boldsymbol{x}^{s}(t,\boldsymbol{b}),\boldsymbol{p}^{s}(t,\boldsymbol{b}^{(0)})=(\boldsymbol{0},-\boldsymbol{b}^{(0)}/t)$ on $t\in[t_{0},+\infty)$. The for loop on lines 2-20 then sequentially applies Propositions 5.1 and 6.1 repeatedly and correctly. We show below that the algorithm converges in at most $m$ iterations by positing the following loop invariant: at iteration $k$, the set $\mathcal{M}^{(k-1)}$ contains at least $k$ indices.
At initialization $k=1$, the equicorrelation set $\mathcal{E}^{(0)}$ is non-empty because $\text{rank}(\bm{A})=m$ by assumption, and so the set $\mathcal{M}^{(0)}$ is non-empty as well, with $|\mathcal{M}^{(0)}|\geqslant 1$. Suppose now that the loop invariant holds at $k\geqslant 1$. We may assume $\boldsymbol{p}^{(k)}\neq\boldsymbol{p}^{(k-1)}$, for otherwise $t^{(k)}=0$ and we are done. Thus at the next iteration $k+1$ we compute the equicorrelation set $\mathcal{E}^{(k)}$ and find, through Proposition 6.1(ii) and (iii), that $\mathcal{E}^{(k)}$ adds at least one index corresponding to a column vector of $\bm{A}$ linearly independent from $\{\bm{A}\boldsymbol{e}_{k}\}_{k\in\mathcal{E}(\boldsymbol{p}^{(k-1)})}$. Hence the subset $\mathcal{M}^{(k)}$ contains one new index, and so $|\mathcal{M}^{(k)}|\geqslant|\mathcal{M}^{(k-1)}|+1\geqslant k+1$.
We claim there is $K\leqslant m$ such that $\hat{\boldsymbol{v}}^{(K-1)}$ solves the NNLS problem (35) exactly, yielding $\boldsymbol{\xi}^{(K-1)}=\boldsymbol{0}$ and $C^{(K-1)}=+\infty$. Indeed, the loop invariant guarantees $|\mathcal{M}^{(k)}|\geqslant k+1$ and $\mathrm{rank}({\bm{A}})=m$; if $K=m$, then $|\mathcal{M}^{(K-1)}|=m$ and the matrix $\bm{A}_{\mathcal{M}^{(K-1)}}$ is invertible, yielding
$\hat{\boldsymbol{v}}^{\text{LSQ}}_{\mathcal{M}(\boldsymbol{p}^{(K-1)})}(t^{(K-1)},\boldsymbol{b}^{(K-1)})=-t^{(K-1)}\bm{D}^{(K-1)}_{\mathcal{M}^{(K-1)}}\bm{A}^{-1}_{\mathcal{M}^{(K-1)}}\boldsymbol{p}^{(K-1)}$
and
$\boldsymbol{\xi}^{(K-1)}$ $=\bm{A}\bm{D}^{(K-1)}\hat{\boldsymbol{v}}^{\text{LSQ}}_{\mathcal{M}(\boldsymbol{p}^{(K-1)})}(t^{(K-1)},\boldsymbol{b}^{(K-1)})+t^{(K-1)}\boldsymbol{p}^{(k-1)}$
$=\bm{A}_{\mathcal{M}(\boldsymbol{p}^{(K-1)})}\bm{D}^{(K-1)}_{\mathcal{M}(\boldsymbol{p}^{(K-1)})}\left(-t^{(K-1)}\bm{D}^{(K-1)}_{\mathcal{M}^{(K-1)}}\bm{A}^{-1}_{\mathcal{M}^{(K-1)}}\boldsymbol{p}^{(K-1)}\right)+t^{(K-1)}\boldsymbol{p}^{(K-1)}$
$=-t^{(K-1)}\boldsymbol{p}^{(K-1)}+t^{(K-1)}\boldsymbol{p}^{(K-1)}=\boldsymbol{0}.$
The subsequent iterations terminate because $\boldsymbol{\xi}^{(K-1)}=\boldsymbol{0}\implies C^{(K-1)}=+\infty\implies t^{(K)}=0$.
After termination, the vectors $(\boldsymbol{x}^{(K)},\boldsymbol{p}^{(K)})$ constitutes a pair of primal and dual solutions to the (modified) basis pursuit problem (48). Formula (49) follows from line 13 and carrying out the sum until $\boldsymbol{b}^{(0)}=\boldsymbol{b}$. Combining (49) with $\bm{A}\boldsymbol{x}^{(K)}=\boldsymbol{b}^{(K)}$ gives
$\boldsymbol{b}$ $=\bm{A}\boldsymbol{x}^{(K)}-\sum_{k=1}^{K}\left(1-t^{(k)}/t^{(k-1)}\right)\bm{A}_{\mathcal{E}(\boldsymbol{p}^{(k-1)})}\bm{D}^{(k-1)}_{\mathcal{E}(\boldsymbol{p}^{(k-1)})}\boldsymbol{w}^{(k-1)}_{\mathcal{E}(\boldsymbol{p}^{(k-1)})}$
$=\bm{A}\left(\boldsymbol{x}^{(K)}-\sum_{k=1}^{K}\left(1-t^{(k)}/t^{(k-1)}\right)\bm{D}^{(k-1)}\boldsymbol{w}^{(k-1)}\right)=\bm{A}\boldsymbol{x}_{f}.$
It follows that $(\boldsymbol{x}_{f},\boldsymbol{p}_{f})$ is a pair of feasible primal and dual solutions to (BP) and (dBP). ∎
Remark 6.1.
A similar reasoning shows Algorithm 3 works when relaxing the assumption to $\boldsymbol{b}\in\text{Im}(\bm{A})$. In this case, the algorithm terminates in at most $\mathrm{rank}({\bm{A}})$ iterations.
Remark 6.2 (Complexity of the algorithm).
Each line in Algorithm 3 can be computed in strongly polynomial time, including, crucially, the calculation of $\hat{\boldsymbol{v}}^{\text{LSQ}}(t^{(k-1)},\boldsymbol{b}^{(k-1)})$ and $\boldsymbol{w}^{(k-1)}$. Since the iterations converge in at most $m$ steps, Algorithm 3 produces a pair of feasible primal and dual solutions to (BP) and (dBP) in strongly polynomial time.
7 Numerical experiments
This section presents some numerical experiments to compare the accuracy and run times of Algorithm 1 with some state-of-the-art algorithms for computing regularization and solution paths to (BPDN). Specifically, we compare below our algorithms with the glmnet software package (glmnet[61, 76, 30, 37], the MATLAB’s native lasso implementation [^3] This is ostensibly an implementation of the glmnet package to MATLAB, except that it has more options, including for better control of the accuracy. (mlasso), and the fast iterative shrinkage thresholding algorithm (fista) with a strong selection rule [6, 32, 62, 53]. For the special case $t=0$, we compare the performance of Algorithm 1 with regularization path starting from $(t^{(0)},\boldsymbol{p}^{(0)})=(\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty},-\boldsymbol{b}/t^{(0)})$, no regularization path and input $\boldsymbol{p}^{(0)}=-\boldsymbol{b}/\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}$, and no regularization path and input provided by Algorithm 3. All experiments below were carried out in MATLAB on an Apple M3 processor.
Implementations and datasets.
All our algorithms require solving a NNLS problem. For this, we use Meyer’s algorithm [47], an active set NNLS algorithm generalizing the Lawson–Hanson algorithm [41] to arbitrary starting active sets, and so particularly suitable for BPDN. For glmnet, we use the code available at https://github.com/junyangq/glmnet-matlab. For fista, we use the variant described in [16, Algorithm 5, Page 197] together with the selection rule described in [62, Equation 7]. For our numerical experiments, we use the benchmark datasets provided by Lorenz et al. [44, Tables 1-2, Section 4], tailored specifically for BPDN and BP and to use for comparing different numerical algorithms. Each dataset comprises a triplet $(\bm{A},\boldsymbol{b},\boldsymbol{x}_{\text{BP}}^{s})$, where $\boldsymbol{x}_{\text{BP}}^{s}=\operatorname*{arg\,min}_{\boldsymbol{x}\in\mathbb{R}^{m}}\left\|{\boldsymbol{x}}\right\|_{1}\ \text{s.t.}\ \bm{A}\boldsymbol{x}=\boldsymbol{b}$. We use in total ten different datasets, corresponding to some of their largest dense matrix ($m=1024,\,n=8192$) with six different observed data and solution vectors and their largest sparse matrix ($m=8192,\,n=49152$) with four different observed data and solution vectors. The solution vectors have either high dynamic range (HDR) or low dynamic range (LDR), and their support either satisfy the so-called Exact Recovery Condition (ERC) [65], an extended form of it (extERC) or, for dense matrices only, no exact recovery condition (noERC) (see [44] for details). All our MATLAB implementations, external software and datasets will be bundled and made publicly available on Zotero.
7.1 Accuracy checks
We compare how accurately glmnet, fista and mlasso compute solutions to (BPDN) and (dBPDN) (via the optimality conditions $t\boldsymbol{p}^{s}(t,\boldsymbol{b})=\bm{A}\boldsymbol{x}^{s}(t,\boldsymbol{b})-\boldsymbol{b}$) across solution paths. We use dataset 548 (dense $1024\times 8192$ matrix, LDR, noERC) and dataset 474 (sparse $8192\times 49512$ matrix, LDR, extERC) from [44], their most computationally demanding datasets. For each dataset, we first compute the entire solution path $t\mapsto(\boldsymbol{x}^{s}(t,\boldsymbol{b}),\boldsymbol{p}^{s}(t,\boldsymbol{b}))$ using Algorithm 2 and identify all $K$ kinks in $t$ in the solution path (see Remark 5.1). Then we run Algorithm 1, glmnet, fista, and mlasso (for dataset 584 as mlasso does not support sparse matrices) at the kinks $\{t^{(k)}\}_{k=0}^{K}$, starting from $t^{(0)}=\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}$.
We perform two runs for each dataset. In the first run, we use the default options of glmnet and mlasso and a relative tolerance of $10^{-4}$ for fista (i.e., the algorithm stops when the relative difference in $\ell_{\infty}$-norm of the dual updates is less than $10^{-4}$). In the second run, we use harsher tolerances ($\text{thresh}=10^{-13}$ for glmnet, $\text{RelTol}=10^{-8}$ for mlasso, and a relative tolerance of $10^{-8}$ for fista). In both runs, we use a tolerance of $10^{-8}$ in Algorithms 1 and 2 when evaluating their equicorrelation sets on line 3. After convergence of each algorithm, we evaluate the dual objective function in (dBPDN) and number of nonzero components of their primal solutions along the solution path.
Figures 1 and 2 below show the relative errors of the dual objective function and the dual solution with respect to Algorithm 1, as well as the number of nonzero components of the primal solutions near the end of the solution path. Algorithm 2 essentially coincides with Algorithm 1 in all cases. More importantly, Algorithm 1 always achieved better optimality in its dual objective function compared to the other algorithms, hence why we use it as the measure for computing all relative errors.
At default tolerance, mlasso and fista performed reasonably well while glmnet performed poorly. In particular, glmnet produced a numerical solution to BPDN with many more nonzero components than all other algorithms. This even remained the case in dataset # 474 with a harsh tolerance: glmnet produced a numerical solution to BPDN with a higher number of nonzero components compared to fista and Algorithm 1. Taken together, we see that Algorithm 1 is the clear winner when it comes to accuracy, with mlasso and fista performing well, and with glmnet performing poorly, even when using harsher tolerances.
Figure 1: Relative error of the dual objective function with respect to Algorithm 1 (first row), relative error in $\ell_{\infty}$-norm of the dual solution with respect to Algorithm 1 (second row), and number of nonzero components of the primal solutions at the end of the solution paths. The dataset used is # 548 from [44]. Left: Default tolerances ($\text{thresh}=10^{-4}$ for glmnet, $\text{RelTol}=10^{-4}$ for mlasso, and a relative tolerance of $10^{-04}$ for fista). Right: Harsher tolerances ($\text{thresh}=10^{-13}$ for glmnet, $\text{RelTol}=10^{-8}$ for mlasso, and a relative tolerance of $10^{-8}$ for fista).
Figure 2: Relative error of the dual objective function with respect to Algorithm 1 (first row), relative error in $\ell_{\infty}$-norm of the dual solution with respect to Algorithm 1 (second row), and number of nonzero components of the primal solutions at the end of the solution path. The dataset used is # 474 from [44]. Left: Default tolerances ($\text{thresh}=10^{-4}$ for glmnet, $\text{RelTol}=10^{-4}$ for mlasso, and a relative tolerance of $10^{-04}$ for fista). Right: Harsher tolerances ($\text{thresh}=10^{-13}$ for glmnet, $\text{RelTol}=10^{-8}$ for mlasso, and a relative tolerance of $10^{-8}$ for fista).
7.2 Run times comparisons
7.2.1 Computing regularization paths to basis pursuit denoising
Table 1 shows the run times of Algorithm 1, glmnet, mlasso, and fista for computing regularization paths. The regularization paths were generated using 512/1024 logarithmically spaced points for the dense/sparse matrices over $t\in\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}\times[10^{-4},1]$. We also included the point $\{0\}$ for Algorithm 1, which the other algorithms cannot handle. Since we wish to compute accurate regularization paths, we used harsh tolerances for each algorithm, just as in Section 7.1 ($\text{thresh}=10^{-13}$ for glmnet, $\text{RelTol}=10^{-08}$ for mlasso, and a relative tolerance of $081e-$ 08 for fista). Table 1 shows that the overall winner is Algorithm 1; it is significantly faster than fista and mlasso, and slightly faster or comparable to glmnet.
Datasets fista mlasso glmnet Algorithm 1
# 147 (d, HDR, ERC) 7.15 1.63 0.87 0.65
# 148 (d, HDR, extERC) 14.46 1.60 0.86 0.69
# 274 (d, HDR, noERC) 88.70 4.16 1.01 1.22
# 421 (d, LDR, ERC) 17.13 1.57 0.88 0.58
# 422 (d, LDR, extERC) 31.69 1.62 0.89 0.60
# 548 (d, LDR, noERC) 524.21 6.64 1.82 1.06
# 199 (s, HDR, ERC) 52.49 N/A 1.62 0.65
# 200 (s, HDR, extERC) 195.86 N/A 1.61 1.07
# 473 (s, LDR, ERC) 479.12 N/A 1.57 0.50
# 474 (s, LDR, extERC) 645.76 N/A 1.67 0.98
Table 1: Timings (in seconds) for Algorithm 1, glmnet, mlasso and fista. Total times for 512/1024 logarithmically spaced points (d: dense/s: sparse), averaged over 5 runs.
7.2.2 Computing optimal solutions to basis pursuit
Table 2 shows the run times of Algorithm 1 with $t=0$ and using as input the feasible dual solution $\boldsymbol{p}_{f}$ obtained from Algorithm 3, Algorithm 1 with $t=0$ and using as input $-\boldsymbol{b}/\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}$, and Algorithm 1 with a regularization path. We used the ten aforementioned benchmark datasets from [44], and the regularization paths for Algorithm 1 were generated using 512/1024 logarithmically spaced points for the dense/sparse matrices over $t\in\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}\times[10^{-4},1]$ and including $\{0\}$. We did not include MATLAB’s native linear programming solver because we found that it was slow and often failed to converge. We also did not include other solvers for (BP), such as AISS [14], SPGL1 [67] and SeDuMi [59], because earlier work by Tendero et al. [60] did similar comparisons with a different version of Algorithm 1 (valid for $t=0$ only) and found it superior in performance and accuracy.
Table 2 shows that all the methods have similar running times. In addition, we’ve verified that all methods correctly computed the solution vectors provided in the benchmark datasets. Algorithm 1 with regularization path is often slower because it used a large number of points. Using Algorithm 3 to get a warm input for Algorithm 1 performs just as well as using Algorithm 1, and in dataset # 274 it was slightly faster.
Datasets Alg. 1 (RP) Alg. 1 ($t=0$) Algs. 1 and 3
# 147 (d, HDR, ERC) 0.65 0.019 0.030
# 148 (d, HDR, extERC) 0.69 0.041 0.047
# 274 (d, HDR, noERC) 1.22 0.72 0.51
# 421 (d, LDR, ERC) 0.58 0.020 0.025
# 422 (d, LDR, extERC) 0.60 0.035 0.044
# 548 (d, LDR, noERC) 1.06 0.58 1.69
# 199 (s, HDR, ERC) 0.65 0.028 0.034
# 200 (s, HDR, extERC) 1.07 1.23 1.20
# 473 (s, LDR, ERC) 0.50 0.028 0.036
# 474 (s, LDR, extERC) 0.98 1.29 1.93
Table 2: Timings (in seconds) for Algorithm 1 with a regularization path, Algorithm 1 with $t=0$ and using $-\boldsymbol{b}/\left\|{\bm{A}^{\top}\boldsymbol{b}}\right\|_{\infty}$ as input, and Algorithm 1 with $t=0$ and using the feasible dual solution $\boldsymbol{p}_{f}$ obtained from Algorithm 3 as input. Total times for 512/1024 logarithmically spaced points and including $\{0\}$ (d: dense/s: sparse), averaged over 5 runs.
8 Conclusion and future work
In this work, we proved that a minimal selection principle from the theory of differential inclusions (Theorem 3.1) enables one to compute an optimal solution of (dBPDN) from the asymptotic limit of the slow system (12). As the results in Section 3 and 4 show, the slow system can be integrated, yielding the slow solution (30). The slow solution converges in finite time and, at convergence, yields the optimal solution to (dBPDN). From it, one can recover an optimal solution to (BPDN). Taken together, these results yielded Algorithm 1. We also presented, in Section 5, a detailed perturbation analysis of the slow system, including its local and global dependence on the hyperparameter and data. The global continuation of the slow solution provided a rigorous, generalized homotopy algorithm for BPDN, as well as Algorithm 3, a greedy algorithm for computing feasible solutions to basis pursuit in strongly polynomial time. Our numerical experiments showed that Algorithm 1 vastly outperforms the state of the arts in accuracy while also achieving the best overall performance, highlighting its key feature that it neither compromises accuracy nor computational efficiency.
While this work focused largely on BPDN, our results provided a novel solution method for solving a broad class of projected dynamical systems. We therefore expect that this method will be relevant to applications involving variational inequalities and projected dynamical systems. In addition, we expect that our results can be adapted to compute exact or approximate solutions to a broader class of convex polyhedral-constrained optimization problems. This includes: (i) variations of BPDN where the $\ell_{1}$ norm is replaced by a polyhedral norm or the function $\boldsymbol{x}\mapsto\left\|{\bm{M}\boldsymbol{x}}\right\|_{1}$ for some appropriate real matrix $\bm{M}$, (ii) other $\ell_{1}$-regularized problems such as logistic regression, Poisson regression, support vector machines and boosting problems, and (iii) extensions to inequality constraints. In particular, as basis pursuit and linear programming are theoretically equivalent problems [64], the results of this work suggest it may be possible to devise algorithms for linear programming based on differential inclusions. These problems will be investigated in future work.
Acknowledgements
GPL would like to thank Prof. Xiaodong Wang for useful discussions, his early support and encouragement on this project, and Prof. Georg Stadler and Dr. David K. A. Mordecai for their support and encouragement. This research is supported by ONR N00014-22-1-2667.
Conflict of interest
The authors declare that they have no conflict of interest.
Appendix A Mathematical background
We list here important definitions and technical results from convex and functional analysis used in this work. For comprehensive references, we refer the reader to [54, 55, 3, 26, 38, 39].
Notation
All vectors and matrices are denoted in bold typeface. Given an $m\times n$ real matrix $\bm{A}$ and a set of indices $\mathcal{E}\subset\{1,\dots,n\}$, we write $\bm{A}_{\mathcal{E}}$ to denote its $m\times|\mathcal{E}|$ submatrix with columns indexed by $\mathcal{E}$ and we write $\bm{A}^{\top}_{\mathcal{E}}$ to denote its transpose. Similarly, given $\boldsymbol{u}\in\mathbb{R}^{n}$, we write $\boldsymbol{u}_{\mathcal{E}}$ to denote its $|\mathcal{E}|$-dimensional subvector indexed by $\mathcal{E}$.
Notation Meaning
$\{\boldsymbol{e}_{1},\dots,\boldsymbol{e}_{n}\}$ The set of $n$ canonical vectors of $\mathbb{R}^{n}$
$\left\langle\boldsymbol{x}_{1},\boldsymbol{x}_{2}\right\rangle$ The Euclidean scalar product of two vectors $\boldsymbol{x}_{1},\boldsymbol{x}_{2}\in\mathbb{R}^{n}$.
$\mathrm{int}{~{}C}$ Interior of a nonempty subset $C$
$\mathrm{ri}~{}C$ Interior of a nonempty subset $C$ relative to the affine hull of $C$
$\chi_{C}$ The characteristic function of a set C:
   $\chi_{C}(\boldsymbol{x})\coloneqq\begin{cases}0,\;&\text{if}\;x\in C\\ +\infty\;&\text{otherwise}\end{cases}$
$\mathrm{dom~{}}f$ The domain of a function $f$: $\mathrm{dom~{}}f\coloneqq\left\{\boldsymbol{x}\in\mathbb{R}^{n}:f(\boldsymbol{x})<+\infty\right\}$
$\mathrm{dom~{}}\partial f$ The set of points $\boldsymbol{x}\in\mathrm{dom~{}}f$ where the subdifferential $\partial f(\boldsymbol{x})\neq\varnothing$
$f^{*}$ Convex conjugate of a function $f$: $f^{*}(\boldsymbol{s})\coloneqq\sup_{\boldsymbol{x}\in\mathbb{R}^{n}}\left\{\left\langle\boldsymbol{s},\boldsymbol{x}\right\rangle-f(\boldsymbol{x})\right\}$
$\mathrm{proj}_{C}(\boldsymbol{x})$ Projection of $\boldsymbol{x}\in\mathbb{R}^{n}$ onto a closed convex set $C\subset\mathbb{R}^{n}$:
   $\mathrm{proj}_{C}(\boldsymbol{x})\coloneqq\operatorname*{arg\,min}_{\boldsymbol{y}\in C}\left\|\boldsymbol{x}-\boldsymbol{y}\right\|_{2}^{2}$
Definitions
Definition A.1 (Convex sets).
A subset $C\subset\mathbb{R}^{n}$ is convex if for every pair $(\boldsymbol{x}_{1},\boldsymbol{x}_{2})\in C\times C$ and every scalar $\lambda\in(0,1)$, the point $\lambda\boldsymbol{x}_{1}+(1-\lambda)\boldsymbol{x}_{2}$ is contained in $C$.
Definition A.2 (Closed convex polyhedra).
A nonempty set $C\subset\mathbb{R}^{n}$ is a closed convex polyhedron if it can be expressed as
$C\coloneqq\left\{\boldsymbol{x}\in\mathbb{R}^{n}:\langle\boldsymbol{u}_{j},\boldsymbol{x}\rangle\leqslant r_{j}\ \text{for every $j\in\{1,\dots,l\}$}\right\},$
where $\{\boldsymbol{u}_{1},\dots,\boldsymbol{u}_{l}\}\subset\mathbb{R}^{n}$ and $\{r_{1},\dots,r_{l}\}\subset\mathbb{R}$.
Definition A.3 (Convex cones).
A nonempty set $K\subset\mathbb{R}^{n}$ is a cone if $\boldsymbol{0}\in K$ and $\lambda\boldsymbol{x}\in K$ for all $\boldsymbol{x}\in K$ and $\lambda>0$. A cone $K$ is convex if it contains the point $\sum_{j=1}^{k}\eta_{j}\boldsymbol{x}_{j}$ whenever $\boldsymbol{x}_{j}\in K$ and $\eta_{j}\geqslant 0$ for $j\in\{1,\dots,k\}$.
Definition A.4 (Conical hulls).
The conical hull of $k$ vectors $\{\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{k}\}\subset\mathbb{R}^{n}$ is defined as the closed convex cone
$\mathrm{cone}\{\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{k}\}\coloneqq\left\{\sum_{j=1}^{k}\eta_{j}\boldsymbol{x}_{j}:\eta_{j}\geqslant 0\,\text{for every}\ j\in\{1,\dots,k\}\right\}.$
Definition A.5 (Polyhedral cones).
A nonempty cone $K\subset\mathbb{R}^{n}$ is polyhedral if it can be expressed as $K=\mathrm{cone}\{\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{k}\}$ for some finite collection of vectors $\{\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{k}\}\subset\mathbb{R}^{n}$.
Definition A.6 (Proper functions).
A function $f$ defined over $\mathbb{R}^{n}$ is proper if its domain
$\mathrm{dom~{}}f\coloneqq\left\{\boldsymbol{x}\in\mathbb{R}^{n}:f(\boldsymbol{x})<+\infty\right\}$
is nonempty and $f(\boldsymbol{x})>-\infty$ for every $\boldsymbol{x}\in\mathrm{dom~{}}f$.
Definition A.7 (Lower semicontinuity).
A proper function $f\colon\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ is lower semicontinuous at $\boldsymbol{x}\in\mathrm{dom~{}}f$ if for every sequence $\left\{\boldsymbol{x}_{k}\right\}_{k=1}^{+\infty}\subset\mathbb{R}^{n}$ converging to $\boldsymbol{x}$,
$\liminf_{k\to+\infty}f(\boldsymbol{x}_{k})\geqslant f(\boldsymbol{x}).$
We say that $f$ is lower semicontinuous if it is lower semicontinuous at every $\boldsymbol{x}\in\mathrm{dom~{}}f$.
Definition A.8 (Convex functions).
A proper function $f\colon\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ is convex if its domain $\mathrm{dom~{}}f$ is convex and if for every pair $(\boldsymbol{x}_{1},\boldsymbol{x}_{2})\in\mathrm{dom~{}}f\times\mathrm{dom~{}}f$ and every $\lambda\in[0,1]$,
$f(\lambda\boldsymbol{x}_{1}+(1-\lambda)\boldsymbol{x}_{2})\leqslant\lambda f(\boldsymbol{x}_{1})+(1-\lambda)f(\boldsymbol{x}_{2}).$
It is strictly convex if the inequality above is strict whenever $\boldsymbol{x}_{1}\neq\boldsymbol{x}_{2}$ and $\lambda\in(0,1)$, and it is $t$-strongly convex with $t>0$ if for every pair $(\boldsymbol{x}_{1},\boldsymbol{x}_{2})\in\mathrm{dom~{}}f\times\mathrm{dom~{}}f$ and every $\lambda\in[0,1]$,
$f(\lambda\boldsymbol{x}_{1}+(1-\lambda)\boldsymbol{x}_{2})\leqslant\lambda f(\boldsymbol{x}_{1})+(1-\lambda)f(\boldsymbol{x}_{2})-\frac{t}{2}\lambda(1-\lambda)\left\|{\boldsymbol{x}_{1}-\boldsymbol{x}_{2}}\right\|_{2}^{2}.$
Definition A.9 (The set $\Gamma_{0}(\mathbb{R}^{n})$).
The space of proper, lower semicontinuous and convex functions over $\mathbb{R}^{n}$ is denoted by $\Gamma_{0}(\mathbb{R}^{n})$.
Definition A.10 (Differentiability).
Let $f$ be a proper function with $\mathrm{int}{~{}\mathrm{dom~{}}f}\neq\emptyset$. The function $f$ is differentiable at $\boldsymbol{x}\in\mathrm{int}{~{}\mathrm{dom~{}}f}$ if there is $\boldsymbol{s}\in\mathbb{R}^{n}$ such that for every $\boldsymbol{d}\in\mathbb{R}^{n}$,
$f^{\prime}(\boldsymbol{x},\boldsymbol{d})=\left\langle\boldsymbol{s},\boldsymbol{y}\right\rangle.$
If $\boldsymbol{s}$ exists, then it is unique, it is called the gradient of $f$ at $\boldsymbol{x}$, and it is written as $\boldsymbol{s}\equiv\nabla f(\boldsymbol{x})$.
Definition A.11 (Subdifferentiability and subgradients).
A function $f\in\Gamma_{0}(\mathbb{R}^{n})$ is subdifferentiable at $\boldsymbol{x}\in\mathbb{R}^{n}$ if there exists $\boldsymbol{s}\in\mathbb{R}^{n}$ such that for every $\boldsymbol{y}\in\mathrm{dom~{}}f$,
$f(\boldsymbol{y})\geqslant f(\boldsymbol{x})+\left\langle\boldsymbol{s},\boldsymbol{y}-\boldsymbol{x}\right\rangle.$ (51)
In this case, $\boldsymbol{s}$ is called a subgradient of the function $f$ at $\boldsymbol{x}$. The set of subgradients at $\boldsymbol{x}\in\mathbb{R}^{n}$ is called the subdifferential of $f$ at $\boldsymbol{x}$ and is denoted by $\partial f(\boldsymbol{x})$. Moreover:
  • (i)
    When nonempty, the subdifferential of $f$ at $\boldsymbol{x}$ is a closed convex set. The set of points $\boldsymbol{x}\in\mathrm{dom~{}}f$ for which $\partial f(\boldsymbol{x})$ is nonempty is denoted by $\mathrm{dom~{}}\partial f$.
  • (ii)
    The function $f$ has a subgradient $\boldsymbol{s}$ at a point $\boldsymbol{x}\in\mathrm{dom~{}}f\iff f^{\prime}(\boldsymbol{x};\boldsymbol{d})\geqslant\langle\boldsymbol{s},\boldsymbol{d}\rangle$ for every $\boldsymbol{d}\in\mathbb{R}^{n}$ [54, Theorem 23.2, page 216].
  • (iii)
    The function $f$ has a unique subgradient at $\boldsymbol{x}$ if and only if $f$ is differentiable at $\boldsymbol{x}$ and $\partial f(\boldsymbol{x})=\{\nabla f(\boldsymbol{x})\}$ [26, Proposition 5.3, page 23].
  • (iv)
    If $f$ is strictly convex, then the inequality in (51) is strict whenever $\boldsymbol{x}\neq\boldsymbol{y}$. If $f$ is $t$-strongly convex with $t>0$, then $f$ is subdifferentiable at $\boldsymbol{x}\in\mathbb{R}^{n}$ if there exists $\boldsymbol{s}\in\mathbb{R}^{n}$ such that for every $\boldsymbol{y}\in\mathrm{dom~{}}f$,
    $f(\boldsymbol{y})\geqslant f(\boldsymbol{x})+\left\langle\boldsymbol{s},\boldsymbol{y}-\boldsymbol{x}\right\rangle+\frac{t}{2}\left\|{\boldsymbol{x}-\boldsymbol{y}}\right\|_{2}^{2}.$ (52)
Definition A.12 (Monotone and maximal monotone mappings).
Let $F$ denote a set-valued mapping from $\mathbb{R}^{n}$ to $\mathbb{R}^{n}$ with graph $\{(\boldsymbol{x},\boldsymbol{v})\in\mathbb{R}^{n}\times\mathbb{R}^{n}\colon\boldsymbol{v}\in F(\boldsymbol{x})\}$. The set-valued mapping $F$ is monotone if for every $\boldsymbol{x}_{1},\boldsymbol{x}_{2}\in\mathbb{R}^{n}$ and every $\boldsymbol{v}_{1}\in F(\boldsymbol{x}_{1})$, $\boldsymbol{v}_{2}\in F(\boldsymbol{x}_{2})$,
$\langle\boldsymbol{v}_{1}-\boldsymbol{v}_{2},\boldsymbol{x}_{1}-\boldsymbol{x}_{2}\rangle\geqslant 0.$ (53)
It is maximal if no other set-valued mapping $\tilde{F}$ contains strictly the graph of $F$.
If $f\in\Gamma_{0}(\mathbb{R}^{n})$, then its subdifferential is a maximal monotone mapping [55, Theorem 12.17]. If also $f$ is $t$-strongly convex with $t>0$, then for every $\boldsymbol{x}_{1},\boldsymbol{x}_{2}\in\mathbb{R}^{n}$ and every $\boldsymbol{v}_{1}\in\partial f(\boldsymbol{x}_{1})$, $\boldsymbol{v}_{2}\in\partial f(\boldsymbol{x}_{2})$ we have the stronger monotone inequality [54, Corollary 31.5.2],
$\langle\boldsymbol{v}_{1}-\boldsymbol{v}_{2},\boldsymbol{x}_{1}-\boldsymbol{x}_{2}\rangle\geqslant t\left\|{\boldsymbol{x}_{1}-\boldsymbol{x}_{2}}\right\|_{2}^{2}.$ (54)
Definition A.13 (Descent directions).
A vector $\boldsymbol{d}\in\mathbb{R}^{n}$ is a descent direction for a proper function $f$ at $\boldsymbol{x}\in\mathrm{dom~{}}f$ if there exists $\tau>0$ such that $\boldsymbol{x}+\tau\boldsymbol{d}\in\mathrm{dom~{}}f$ and $f(\boldsymbol{x}+\tau\boldsymbol{d})<f(\boldsymbol{x})$.
Definition A.14 (Convex conjugates).
Let $f\in\Gamma_{0}(\mathbb{R}^{n})$. The convex conjugate $f^{*}\colon\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ of $f$ is defined by
$f^{*}(\boldsymbol{s})\coloneqq\sup_{\boldsymbol{x}\in\mathbb{R}^{n}}\left\{\left\langle\boldsymbol{s},\boldsymbol{x}\right\rangle-f(\boldsymbol{x})\right\}.$
By definition, the function $f^{*}$ is in $\Gamma_{0}(\mathbb{R}^{n})$ [26, Definition 4.1, page 17].
Definition A.15 (Coercive functions).
A proper function $f\colon\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ is coercive if for every sequence $\{\boldsymbol{x}_{k}\}_{k=1}^{+\infty}\subset\mathbb{R}^{n}$ with $\lim_{k\to+\infty}\left\|{\boldsymbol{x}_{k}}\right\|_{2}=+\infty$,
$\lim_{k\to+\infty}f(\boldsymbol{x}_{k})=+\infty.$
If $f\in\Gamma_{0}(\mathbb{R}^{n})$, then $f$ is coercive if and only if $\boldsymbol{0}\in\mathrm{ri}\,\mathrm{dom~{}}f^{*}$ [55, Theorem 11.8, page 479].
Definition A.16 (Characteristic functions).
The characteristic function of a nonempty subset $C\subset\mathbb{R}^{n}$ is defined as
$\chi_{C}(\boldsymbol{x})=\begin{cases}0,\;&\text{if}\;x\in C\\ +\infty\;&\text{otherwise}\end{cases}.$
We have $\chi_{C}\in\Gamma_{0}(\mathbb{R}^{n})$ if and only if $C$ is nonempty, closed and convex, and it is coercive if and only if $C$ is bounded.
Definition A.17 (Euclidean projection).
Let $C$ be a nonempty, closed and convex subset of $\mathbb{R}^{n}$ and let $\boldsymbol{x}\in\mathbb{R}^{n}$. We denote by $\mathrm{proj}_{C}(\boldsymbol{x})$ the projection of $\boldsymbol{x}$ on $C$:
$\mathrm{proj}_{C}(\boldsymbol{x})\coloneqq\operatorname*{arg\,min}_{\boldsymbol{y}\in C}\left\|{\boldsymbol{x}-\boldsymbol{y}}\right\|_{2}.$
The projection of $\boldsymbol{x}$ on $C$ exists, is unique, and satisfies the characterization
$\langle\boldsymbol{x}-\mathrm{proj}_{C}(\boldsymbol{x}),\boldsymbol{y}-\mathrm{proj}_{C}(\boldsymbol{x})\rangle\leqslant 0\;\text{for every }\boldsymbol{y}\in C.$ (55)
See [38, Section III, 3.1] for details.
Definition A.18 (Normal cones).
Let $C$ be a nonempty closed convex set. The subdifferential of the characteristic function $\chi_{C}$ at $\boldsymbol{x}\in C$ is the set of normal vectors
$\partial\chi_{C}(\boldsymbol{x})\coloneqq N_{C}(\boldsymbol{x})=\left\{\boldsymbol{s}\in\mathbb{R}^{n}:\langle\boldsymbol{s},\boldsymbol{x}-\boldsymbol{y}\rangle\geqslant 0\ \text{for every $\boldsymbol{y}\in C$}\right\}.$
The set $N_{C}(\boldsymbol{x})$ is called the normal cone of $C$ at $\boldsymbol{x}\in C$, and it is a closed convex cone.
Technical results
Theorem A.1 (The generalized Fermat’s rule).
Let $f\in\Gamma_{0}(\mathbb{R}^{n})$. Then $f$ has a global minimum at $\boldsymbol{x}^{s}$ if and only if $\boldsymbol{0}\in\partial f(\boldsymbol{x}^{s})$.
Proof.
See [55, Theorem 10.1, page 422] for the proof. ∎
Theorem A.2 (Fenchel’s inequality).
Let $f\in\Gamma_{0}(\mathbb{R}^{n})$. For every $\boldsymbol{x},\boldsymbol{s}\in\mathbb{R}^{n}$, the function $f$ and its convex conjugate $f^{*}$ satisfy Fenchel’s inequality:
$f(\boldsymbol{x})+f^{*}(\boldsymbol{s})\geqslant\left\langle\boldsymbol{s},\boldsymbol{x}\right\rangle,$ (56)
with equality if and only if $\boldsymbol{s}\in\partial f(\boldsymbol{x})$, if and only if $\boldsymbol{x}\in\partial f^{*}(\boldsymbol{s})$.
Proof.
See [39, Corollary 1.4.4, page 48]. ∎
Proposition A.1 (The subdifferential of the characteristic function of a closed convex polyhedron).
Let $C\subset\mathbb{R}^{n}$ denote the nonempty, closed convex polyhedron
$C\coloneqq\left\{\boldsymbol{x}\in\mathbb{R}^{n}:\langle\boldsymbol{u}_{j},\boldsymbol{x}\rangle\leqslant r_{j}\ \text{for every $j\in\{1,\dots,l\}$}\right\},$
where $\{\boldsymbol{u}_{1},\dots,\boldsymbol{u}_{l}\}\subset\mathbb{R}^{n}$ and $\{r_{1},\dots,r_{l}\}\subset\mathbb{R}$. Let $\mathcal{E}(\boldsymbol{x})=\{j\in\{1,\dots,l\}:\langle\boldsymbol{u}_{j},\boldsymbol{x}\rangle=r_{j}\}$ denote the set of active constraints at $\boldsymbol{x}\in\mathbb{R}^{n}$. The subdifferential of the characteristic function of $C$ at $\boldsymbol{x}$ is the polyhedral cone
$\partial\chi_{C}(\boldsymbol{x})=\mathrm{cone}\{\boldsymbol{u}_{j}:j\in\mathcal{E}(\boldsymbol{x})\}=\left\{\sum_{j\in\mathcal{E}(\boldsymbol{x})}\eta_{j}\boldsymbol{u}_{j}:\eta_{j}\geqslant 0\right\}.$
Proof.
This follows from Definitions A.2 and A.18. See [38, Example 5.2.6] for details. ∎
Proposition A.2 (Properties of subdifferentials and some calculus rules).
  • (i)
    Let $f_{1},f_{2}\in\Gamma_{0}(\mathbb{R}^{n})$ and assume $\mathrm{ri}\,\mathrm{dom~{}}f_{1}\cap\mathrm{ri}\,\mathrm{dom~{}}f_{2}\neq\emptyset$. Then $f_{1}+f_{2}\in\Gamma_{0}(\mathbb{R}^{n})$ and for any $\boldsymbol{x}\in\mathrm{dom~{}}f_{1}\cap\mathrm{dom~{}}f_{2}$, we have $\partial(f_{1}+f_{2})(\boldsymbol{x})=\partial f_{1}(\boldsymbol{x})+\partial f_{2}(\boldsymbol{x})$.
  • (ii)
    Let $f_{1},f_{2}\in\Gamma_{0}(\mathbb{R}^{n})$ and assume $\mathrm{ri}\,\mathrm{dom~{}}f_{1}\cap\mathrm{ri}\,\mathrm{dom~{}}f_{2}\neq\emptyset$. Then for any $\boldsymbol{s}\in\mathrm{dom~{}}f_{1}^{*}+\mathrm{dom~{}}f_{2}^{*}$, we have $(f_{1}+f_{2})^{*}(\boldsymbol{s})=\inf_{\boldsymbol{s}_{1}+\boldsymbol{s}_{2}=\boldsymbol{s}}\left\{f_{1}^{*}(\boldsymbol{s}_{1})+f_{2}^{*}(\boldsymbol{s}_{2})\right\}$.
  • (iii)
    Let $f\in\Gamma_{0}(\mathbb{R}^{n})$, $g\in\Gamma_{0}(\mathbb{R}^{m})$ and assume $\mathrm{Im}({\bm{A}})\cap\mathrm{ri}\,\mathrm{dom~{}}g\neq\emptyset$. Then the function $\boldsymbol{x}\mapsto g(\bm{A}\boldsymbol{x})$ is in $\Gamma_{0}(\mathbb{R}^{n})$, $\mathrm{dom~{}}(g\circ\bm{A})^{*}=\bm{A}^{\top}\mathrm{dom~{}}g^{*}$, and for all $\boldsymbol{x}\in\mathbb{R}^{n}$ satisfying $\bm{A}\boldsymbol{x}\in\mathrm{dom~{}}g$ we have $\partial(g\circ\bm{A})(\boldsymbol{x})=\bm{A}^{\top}\partial g(\bm{A}\boldsymbol{x})$.
  • (iv)
    Let $f\in\Gamma_{0}(\mathbb{R}^{n})$, $g\in\Gamma_{0}(\mathbb{R}^{m})$ and assume $\mathrm{ri}\,\mathrm{dom~{}}f\cap\mathrm{ri}\,\mathrm{dom~{}}(g\circ\bm{A})\neq\emptyset$. Then for any $\boldsymbol{x}\in\mathrm{dom~{}}f\cap\mathrm{dom~{}}(g\circ\bm{A})$, we have $\partial(f+g\circ\bm{A})=\partial f(\boldsymbol{x})+\bm{A}^{\top}\partial g(\bm{A}\boldsymbol{x})$.
Proof.
See [39, Corollary 3.1.2] for the proof of (i), [39, Theorem 2.3.3] for the proof of (ii), and [39, Theorem 2.2.1, Theorem 2.2.3, and Theorem 3.2.1.] for the proof of (iii). The proof of (iv) follows by using (i) with $f_{1}=f$ and $f_{2}=g\circ\bm{A}$ and then using (iii). ∎
Theorem A.3 (The primal problem and the dual problem).
Let $f_{1}\in\Gamma_{0}(\mathbb{R}^{n})$, $f_{2}\in\Gamma_{0}(\mathbb{R}^{m})$ and let $\bm{A}$ denote a real $m\times n$ matrix. Consider the “primal” minimization problem
$\inf_{\boldsymbol{x}\in\mathbb{R}^{n}}\left\{f_{1}(\boldsymbol{x})+f_{2}(\bm{A}\boldsymbol{x})\right\}.$ (57)
Assume
$\boldsymbol{0}\in\mathrm{ri}(\bm{A}~{}\mathrm{dom~{}}f_{1}-\mathrm{dom~{}}f_{2})\quad\text{and}\quad\boldsymbol{0}\in\mathrm{ri}(\mathrm{dom~{}}f_{1}^{*}+\bm{A}^{\top}\mathrm{dom~{}}f_{2}^{*}).$ (58)
Then:
  • (i)
    The primal problem (57) has at least one solution. If $\boldsymbol{x}^{s}$ denote such a solution, then it satisfies the first-order optimality condition
    $\boldsymbol{0}\in\partial(f_{1}+f_{2}\circ\bm{A})(\boldsymbol{x}^{s})\quad\iff\quad\boldsymbol{0}\in\partial f_{1}(\boldsymbol{x}^{s})+\bm{A}^{\top}\partial f_{2}(\bm{A}\boldsymbol{x}^{s}).$ (59)
  • (ii)
    The “dual” maximization problem
    $\sup_{\boldsymbol{p}\in\mathbb{R}^{m}}\left\{-f_{1}^{*}(-\bm{A}^{*}\boldsymbol{p})-f_{2}^{*}(\boldsymbol{p})\right\}$ (60)
    has at least one solution and is equal in value to the primal problem (57). If $\boldsymbol{p}^{s}$ denote such a solution, then it satisfies the first-order optimality condition
    $\boldsymbol{0}\in\partial(f_{1}^{*}\circ(-\bm{A}^{\top})+f_{2}^{*})(\boldsymbol{p}^{s})\quad\iff\quad\boldsymbol{0}\in-\bm{A}\partial f_{1}^{*}(-\bm{A}^{\top}\boldsymbol{p}^{s})+\partial f_{2}^{*}(\boldsymbol{p}^{s}).$ (61)
  • (iii)
    The optimality conditions (59) and (61) are equivalent and can be written as
    $\boldsymbol{p}^{s}\in\partial f_{2}(\bm{A}\boldsymbol{x}^{s})\ \text{and}\ -\bm{A}^{\top}\boldsymbol{p}^{s}\in\partial f_{1}(\boldsymbol{x}^{s}).$ (62)
Proof.
See [54, Theorem 31.2 and Corollary 31.2.1] and [3, Proposition 1, Theorem 1 and Theorem 2, pages 163–167] for a proof. The equivalences in (59) and (61) follow from (58) and the rules in Proposition A.2. The equivalence in (62) follows from Fenchel’s inequality (56). ∎
Remark A.1.
The assumptions in (58) read: there exists $\boldsymbol{x}\in\mathrm{dom~{}}f_{1}$ such that $\bm{A}\boldsymbol{x}\in\mathrm{ri}\ \mathrm{dom~{}}f_{2}$ and there exists $\boldsymbol{p}\in\mathrm{dom~{}}f_{2}^{*}$ such that $-\bm{A}^{\top}\boldsymbol{p}\in\mathrm{dom~{}}f_{1}^{*}$. If the first assumption holds, then Proposition A.2 implies the function $\boldsymbol{x}\mapsto f_{1}(\boldsymbol{x})+f_{2}(\bm{A}\boldsymbol{x})$ is in $\Gamma_{0}(\mathbb{R}^{n})$. Moreover,
$\mathrm{ri}\,\mathrm{dom~{}}(f_{1}+(f_{2}\circ\bm{A}))^{*}=\mathrm{ri}(\mathrm{dom~{}}f_{1}^{*}+\mathrm{dom~{}}(f_{2}\circ\bm{A})^{*})=\mathrm{ri}(\mathrm{dom~{}}f_{1}^{*}+\bm{A}^{\top}\mathrm{dom~{}}f_{2}^{*}).$
In particular, the second assumption becomes equivalent to coercivity of the function $\boldsymbol{x}\mapsto f_{1}(\boldsymbol{x})+f_{2}(\bm{A}\boldsymbol{x})$ (see Definition A.15).
Appendix B Detailed Proofs
B.1 Proof of Proposition 3.1
Part 1. First, we show there is $\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})>0$ such that $\left\|{-\bm{A}^{\top}(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}))}\right\|_{\infty}\leqslant 1$ for every $\Delta\in[0,\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})]$. Multiply the vector $-\bm{A}^{\top}(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$ by the matrix of signs $\bm{D}(\boldsymbol{p}_{0})$ and take the inner product with respect to the unit vector $\boldsymbol{e}_{j}$ with $j\in\{1,\dots,n\}$:
$\langle-\bm{D}(\boldsymbol{p}_{0})\bm{A}^{\top}(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})),\boldsymbol{e}_{j}\rangle=\langle-\bm{D}(\boldsymbol{p}_{0})\bm{A}^{\top}\boldsymbol{p}_{0},\boldsymbol{e}_{j}\rangle-\Delta\langle\bm{D}(\boldsymbol{p}_{0})\bm{A}^{\top}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\boldsymbol{e}_{j}\rangle.$ (63)
By definition of the equicorrelation set (4) and the matrix of signs $\bm{D}(\boldsymbol{p}_{0})$, we have
$0\leqslant\langle-\bm{D}(\boldsymbol{p}_{0})\bm{A}^{\top}\boldsymbol{p}_{0},\boldsymbol{e}_{j}\rangle\leqslant 1\ \text{for every $j\in\{1,\dots,n\}$.}$
We now proceed according to whether $j\in\mathcal{E}(\boldsymbol{p}_{0})$ or $j\in\mathcal{E}^{\mathsf{C}}(\boldsymbol{p}_{0})$.
First, suppose $j\in\mathcal{E}(\boldsymbol{p}_{0})$. Then $\langle-\bm{D}(\boldsymbol{p}_{0})\bm{A}^{\top}\boldsymbol{p}_{0},\boldsymbol{e}_{j}\rangle=1$ and $\langle\bm{D}(\boldsymbol{p}_{0})\bm{A}^{\top}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\boldsymbol{e}_{j}\rangle\geqslant 0$ by the KKT condition (17a). If the latter is strictly positive, then we can increase $\Delta$ until the left-hand-side of (63) is equal to $-1$. The smallest such number is
$\Delta_{\mathcal{E}(\boldsymbol{p}_{0})}=\min_{j\in\mathcal{E}(\boldsymbol{p}_{0})}2/\langle\bm{D}(\boldsymbol{p}_{0})\bm{A}^{\top}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\boldsymbol{e}_{j}\rangle,$
[…]
Next, suppose $j\in\mathcal{E}^{\mathsf{C}}(\boldsymbol{p}_{0})$ and $\langle\bm{D}(\boldsymbol{p}_{0})\bm{A}^{\top}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\boldsymbol{e}_{j}\rangle\geqslant 0$. The same reasoning as above shows
$\Delta_{\mathcal{E}^{\mathsf{C}}(\boldsymbol{p}_{0}),\geqslant}=\min_{\begin{subarray}{c}j\in\mathcal{E}^{\mathsf{C}}(\boldsymbol{p}_{0})\\ \langle\bm{D}(\boldsymbol{p}_{0})\bm{A}^{\top}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\boldsymbol{e}_{j}\rangle\geqslant 0\end{subarray}}\frac{1-\langle\bm{D}(\boldsymbol{p}_{0})\bm{A}^{\top}\boldsymbol{p}_{0},\boldsymbol{e}_{j}\rangle}{\langle\bm{D}(\boldsymbol{p}_{0})\bm{A}^{\top}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\boldsymbol{e}_{j}\rangle}$
[…]
Finally, suppose $j\in\mathcal{E}^{\mathsf{C}}(\boldsymbol{p}_{0})$ and $\langle\bm{D}(\boldsymbol{p}_{0})\bm{A}^{\top}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\boldsymbol{e}_{j}\rangle<0$. Then we can increase $\Delta$ until the left-hand-side of (63) is equal to $1$. The smallest such number is
$\Delta_{\mathcal{E}^{\mathsf{C}}(\boldsymbol{p}_{0}),<}=\min_{\begin{subarray}{c}j\in\mathcal{E}^{\mathsf{C}}(\boldsymbol{p}_{0})\\ \langle\bm{D}(\boldsymbol{p}_{0})\bm{A}^{\top}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\boldsymbol{e}_{j}\rangle<0\end{subarray}}\frac{1+\langle\bm{D}(\boldsymbol{p}_{0})\bm{A}^{\top}\boldsymbol{p}_{0},\boldsymbol{e}_{j}\rangle}{|\langle\bm{D}(\boldsymbol{p}_{0})\bm{A}^{\top}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\boldsymbol{e}_{j}\rangle|}.$
Hence we find $\left\|{-\bm{A}^{\top}(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}))}\right\|_{\infty}\leqslant 1$ for all $\Delta\in[0,\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})]$, where
$\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ $\coloneqq\min{(\Delta_{\mathcal{E}(\boldsymbol{p}_{0})},\Delta_{\mathcal{E}^{\mathsf{C}}(\boldsymbol{p}_{0}),\geqslant},\Delta_{\mathcal{E}^{\mathsf{C}}(\boldsymbol{p}_{0}),<})}$
$\equiv\min_{j\in\{1,\dots,n\}}\left\{\frac{\mathrm{sgn}{\left\langle\bm{D}(\boldsymbol{p}_{0})\bm{A}^{\top}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\boldsymbol{e}_{j}\right\rangle}-\left\langle\bm{D}(\boldsymbol{p}_{0})\bm{A}^{\top}\boldsymbol{p}_{0},\boldsymbol{e}_{j}\right\rangle)}{\left\langle\bm{D}(\boldsymbol{p}_{0})\bm{A}^{\top}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\boldsymbol{e}_{j}\right\rangle}\right\}.$
Now, for the equivalence, note the assumption $\mathrm{rank}({\bm{A}})=m$ implies $\bm{A}^{\top}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})=\boldsymbol{0}\iff\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})=\boldsymbol{0}$. Clearly $\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})=\boldsymbol{0}$ implies $\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})=+\infty$. Finally, $\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})=+\infty$ only if $\bm{A}^{\top}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})=\boldsymbol{0}$, which is equivalent to $\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})=\boldsymbol{0}$.
Part 2. Using Part 1, we find $V(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b});t,\boldsymbol{b})<+\infty$ for every $\Delta\in[0,\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})]$. In particular, we can write
$V(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b});t,\boldsymbol{b})$ $=\frac{t}{2}\left\|{\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})}\right\|_{2}^{2}+\langle\boldsymbol{b},\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\rangle$
$=t\Delta\langle\boldsymbol{p}_{0},\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\rangle+\Delta^{2}t\left\|{\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})}\right\|_{2}^{2}/2$
$\qquad\qquad+\Delta\langle\boldsymbol{b},\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\rangle+\frac{t}{2}\left\|{\boldsymbol{p}_{0}}\right\|_{2}^{2}+\langle\boldsymbol{b},\boldsymbol{p}_{0}\rangle$
$=t\Delta\langle\boldsymbol{p}_{0},\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\rangle+\Delta^{2}t\left\|{\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})}\right\|_{2}^{2}/2$
$\qquad\qquad+\Delta\langle\boldsymbol{b},\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\rangle+V(\boldsymbol{p}_{0};t,\boldsymbol{b}).$
Substituting (17b) in the above and rearranging yields
$V(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b});t,\boldsymbol{b})-V(\boldsymbol{p}_{0};t,\boldsymbol{b})=\Delta(t\Delta/2-1)\left\|{\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})}\right\|_{2}^{2}.$
This proves Equation (19). Finally, suppose $\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\neq\boldsymbol{0}$, meaning $0<\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})<+\infty$. Taking $\Delta\in(0,\min(\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}),2/t))$ in (19) yields $V(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b});t,\boldsymbol{b})-V(\boldsymbol{p}_{0};t,\boldsymbol{b})<0$. Hence $\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ is a descent direction.
Part 3. Let $\Delta\in[0,\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})]$ and $\boldsymbol{q}\in\mathbb{R}^{m}$. Since $-\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\in\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0};t,\boldsymbol{b})$ and $V(\cdot;t,\boldsymbol{b})$ is $t$-strongly convex, the subdifferentiability property implies
$V(\boldsymbol{q};t,\boldsymbol{b})$ $\geqslant V(\boldsymbol{p}_{0};t,\boldsymbol{b})+\langle-\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\boldsymbol{q}-\boldsymbol{p}_{0}\rangle+\frac{t}{2}\left\|{\boldsymbol{q}-\boldsymbol{p}_{0}}\right\|_{2}^{2}$
$=V(\boldsymbol{p}_{0};t,\boldsymbol{b})+\langle-\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\boldsymbol{q}-(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}))\rangle+\langle-\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\rangle$
$\qquad+\frac{t}{2}\left\|{\boldsymbol{q}-(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}))+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})}\right\|_{2}^{2}$
$=V(\boldsymbol{p}_{0};t,\boldsymbol{b})+\langle-\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\boldsymbol{q}-(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}))\rangle-\Delta\left\|{\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})}\right\|_{2}^{2}$
$\qquad+\frac{t}{2}\left\|{\boldsymbol{q}-(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}))}\right\|_{2}^{2}+\frac{\Delta^{2}t}{2}\left\|{\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})}\right\|_{2}^{2}$
$\qquad\qquad+t\Delta\langle\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\boldsymbol{q}-(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}))\rangle.$
Use Equation (19) in the inequality above to simplify the right hand side:
$V(\boldsymbol{q};t,\boldsymbol{b})$ $\geqslant V(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b});t,\boldsymbol{b})+\langle-(1-t\Delta)\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\boldsymbol{q}-(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}))\rangle$
$\qquad\qquad+\frac{t}{2}\left\|{\boldsymbol{q}-(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}))}\right\|_{2}^{2}.$
By definition of subdifferentiability, this means
$-(1-t\Delta)\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\in\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b});t,\boldsymbol{b})\ \text{for every $\Delta\in[0,\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})]$}.$
Part 4. The inclusions and identity follow for $\Delta=0$, so suppose $\Delta\in(0,\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$. Assume $j\notin\mathcal{E}(\boldsymbol{p}_{0})$. By construction from Part 1, we have $j\notin\mathcal{E}(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$, hence the contrapositive $\mathcal{E}(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}))\subset\mathcal{E}(\boldsymbol{p}_{0})$.
Now, let $\boldsymbol{s}\in\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b});t,\boldsymbol{b})$. Then there is $\hat{\boldsymbol{u}}\in\mathbb{R}^{n}$ with $\hat{\boldsymbol{u}}_{\mathcal{E}(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0},t))}\geqslant\boldsymbol{0}$ and $\hat{\boldsymbol{u}}_{\mathcal{E}^{\mathsf{C}}(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0},t))}=\boldsymbol{0}$ such that
$\boldsymbol{s}=\boldsymbol{b}+t(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}))-\bm{A}_{\mathcal{E}(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0},t))}\bm{D}_{\mathcal{E}(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0},t))}\hat{\boldsymbol{u}}_{\mathcal{E}(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0},t))}.$
In particular,
$\boldsymbol{s}-t\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})=\boldsymbol{b}+t\boldsymbol{p}_{0}-\bm{A}_{\mathcal{E}(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0},t))}\bm{D}_{\mathcal{E}(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0},t))}\hat{\boldsymbol{u}}_{\mathcal{E}(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0},t))}.$
From the inclusion $\mathcal{E}(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}))\subset\mathcal{E}(\boldsymbol{p}_{0})$, we find $\boldsymbol{s}\in\{t\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\}+\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0};t,\boldsymbol{b})$. Since $\boldsymbol{s}$ was arbitrary, we deduce the inclusion
$\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b});t,\boldsymbol{b})\subset\{t\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\}+\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0};t,\boldsymbol{b}).$
Part 5. Next, we prove identity (20). Let $\boldsymbol{s}=\mathrm{proj}_{\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b});t,\boldsymbol{b})}(\boldsymbol{0})$ and use both the previous inclusion and subdifferentiability to find
$V(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b});t,\boldsymbol{b})$ $\geqslant V(\boldsymbol{p}_{0};t,\boldsymbol{b})+\langle\boldsymbol{s}-t\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})-\boldsymbol{p}_{0}\rangle$
$\qquad\qquad+\frac{t}{2}\left\|{\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})-\boldsymbol{p}_{0}}\right\|_{2}^{2}$
$=V(\boldsymbol{p}_{0};t,\boldsymbol{b})+\langle\boldsymbol{s}-t\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\rangle$
$\qquad\qquad+\frac{\Delta^{2}t}{2}\left\|{\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})}\right\|_{2}^{2}$
$=V(\boldsymbol{p}_{0};t,\boldsymbol{b})+\Delta\langle\boldsymbol{s},\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\rangle-\frac{\Delta^{2}t}{2}\left\|{\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})}\right\|_{2}^{2}$
Using Equation (19) in the previous inequality and simplifying yields
$(t\Delta-1)\left\|{\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})}\right\|_{2}^{2}\geqslant\langle\boldsymbol{s},\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\rangle.$ (64)
Next, we use the inclusion $\mathcal{E}(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}))\subset\mathcal{E}(\boldsymbol{p}_{0})$, which was proven in Part 4, and subdifferentiability to find
$V(\boldsymbol{p}_{0};t,\boldsymbol{b})$ $\geqslant V(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b});t,\boldsymbol{b})+\langle\boldsymbol{s},\boldsymbol{p}_{0}-(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}))\rangle$
$\qquad\qquad+\frac{t}{2}\left\|{\boldsymbol{p}_{0}-(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}))}\right\|_{2}^{2}$
$=V(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b});t,\boldsymbol{b})-\Delta\langle\boldsymbol{s},\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\rangle+\frac{\Delta^{2}t}{2}\left\|{\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})}\right\|_{2}^{2}.$
Using Equation (19) and substituting in the above yields
$\langle\boldsymbol{s},\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})\rangle\geqslant(t\Delta-1)\left\|{\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})}\right\|_{2}^{2},$ (65)
Combining inequalities (64) and (65) yields the equality
$\langle\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),(1-t\Delta)\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})+\boldsymbol{s}\rangle=0.$ (66)
Finally, we use the projection characterization (55) with
$\boldsymbol{x}=\boldsymbol{0},\,C=\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b});t,\boldsymbol{b})\ \text{and}\ \boldsymbol{y}=-(1-t\Delta)\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}),$
and use (66) to get the inequality
$\langle\boldsymbol{s},(1-t\Delta)\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})+\boldsymbol{s}\rangle\leqslant 0.$ (67)
However, multiplying (66) by $(1-t\Delta)$ and adding it to (67) yields
$\langle(1-t\Delta)\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})+\boldsymbol{s},(1-t\Delta)\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})+\boldsymbol{s}\rangle=\left\|{(1-t\Delta)\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})+\boldsymbol{s}}\right\|_{2}^{2}\leqslant 0.$
We deduce $\boldsymbol{s}=-(1-t\Delta)\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})$, that is,
$-\mathrm{proj}_{\partial_{\boldsymbol{p}}V(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b});t,\boldsymbol{b})}(\boldsymbol{0})\equiv\boldsymbol{d}(\boldsymbol{p}_{0}+\Delta\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b});t,\boldsymbol{b})=(1-t\Delta)\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}).$
B.2 Proof of Lemma 4.1
Suppose $\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})<+\infty$. By Proposition 3.1(iii) and Theorem 4.1, we have
$\frac{d\boldsymbol{p}}{d\tau}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})),\frac{d\boldsymbol{p}_{a}}{d\tau}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))\in-\partial_{\boldsymbol{p}}V(\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}));t,\boldsymbol{b}).$
As a consequence, we can extend the analysis done in Theorem 4.1 to the value $\tau=\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$. That is, $\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))=\boldsymbol{p}_{a}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$, and so (23) holds at $\tau=\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$.
Next, we prove inequality (28). It suffices to show
$\boldsymbol{d}(\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}));t,\boldsymbol{b})\neq e^{-t\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}).$
Indeed, the projection onto the closed and convex set $\partial_{\boldsymbol{p}}V(\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}));t,\boldsymbol{b})$ is unique (see A.17), so the subgradients are different if and only if their norms are different. First, we apply Theorem 4.1 with the initial value $\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))$ at $\tau=\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ to find
$\boldsymbol{p}(\tau)=\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}))+f(\tau-\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}),t)\boldsymbol{d}(\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}));t,\boldsymbol{b})$ (68)
in the time interval $[\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\tau_{*}(\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}));t,\boldsymbol{b})]$. Assume for a contradiction that $\boldsymbol{d}(\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}));t,\boldsymbol{b})=e^{-t\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})$. Then (68) can be developed further into
$\boldsymbol{p}(\tau)=\boldsymbol{p}_{0}+\left(\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})+f(\tau-\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}),t)e^{-t\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})}\right)\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b}).$
However, this would imply
$\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})+f(\tau-\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}),t)e^{-t\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})}>\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$
for every $\tau\in(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}),\tau_{*}(\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}));t,\boldsymbol{b})]$, contradicting the definition of the number $\Delta_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})$ in (18). We conclude $\boldsymbol{d}(\boldsymbol{p}(\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b}));t,\boldsymbol{b})\neq e^{-t\tau_{*}(\boldsymbol{p}_{0};t,\boldsymbol{b})}\boldsymbol{d}(\boldsymbol{p}_{0};t,\boldsymbol{b})$.
B.3 Proof of Lemma 5.1
Part 1. Use Proposition 3.1 with hyperparameter $t_{0}+\delta_{0}$ and data $\boldsymbol{b}+\boldsymbol{\delta}\boldsymbol{b}$ to get
$\boldsymbol{d}(\boldsymbol{p}_{0};t_{0}+\delta_{0},\boldsymbol{b}+\boldsymbol{\delta}\boldsymbol{b})=\bm{A}\bm{D}(\boldsymbol{p}_{0})\hat{\boldsymbol{u}}(\boldsymbol{p}_{0};t_{0}+\delta_{0},\boldsymbol{b}+\boldsymbol{\delta}\boldsymbol{b})-(\boldsymbol{b}+\boldsymbol{\delta}\boldsymbol{b})-(t_{0}+\delta_{0})\boldsymbol{p}_{0},$
where
$\hat{\boldsymbol{u}}(\boldsymbol{p}_{0};t_{0}+\delta_{0},\boldsymbol{b}+\boldsymbol{\delta}\boldsymbol{b})\in\operatorname*{arg\,min}_{\boldsymbol{u}\in\mathbb{R}^{n}}\left\|{\bm{A}\bm{D}(\boldsymbol{p}_{0})\boldsymbol{u}-\boldsymbol{b}-\boldsymbol{\delta}\boldsymbol{b}-(t_{0}+\delta_{0})\boldsymbol{p}_{0}}\right\|_{2}^{2}$
subject $\qquad\qquad\text{subject to}\ \begin{cases}\boldsymbol{u}_{\mathcal{E}(\boldsymbol{p}_{0})}\geqslant\boldsymbol{0}\\ \boldsymbol{u}_{\mathcal{E}^{\mathsf{C}}(\boldsymbol{p}_{0})}=\boldsymbol{0}.\end{cases}$
Now let $\boldsymbol{v}\in\mathbb{R}^{m}$ and use the change of variables $\boldsymbol{u}=\hat{\boldsymbol{u}}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})+\boldsymbol{v}$ in the optimization problem above. The constraints of the problem become $\boldsymbol{v}_{\mathcal{E}(\boldsymbol{p}_{0})}\geqslant-\hat{\boldsymbol{u}}_{\mathcal{E}(\boldsymbol{p}_{0})}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})$ and $\boldsymbol{v}_{\mathcal{E}^{\mathsf{C}}(\boldsymbol{p}_{0})}=\boldsymbol{0}$, while the objective function becomes
$\left\|{\bm{A}\bm{D}(\boldsymbol{p}_{0})\boldsymbol{u}-\boldsymbol{b}-\boldsymbol{\delta}\boldsymbol{b}-(t_{0}+\delta_{0})\boldsymbol{p}_{0}}\right\|_{2}^{2}$
$\qquad\qquad\qquad=\left\|{\bm{A}\bm{D}(\boldsymbol{p}_{0})\hat{\boldsymbol{u}}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})+\bm{A}\bm{D}(\boldsymbol{p}_{0})\boldsymbol{v}-\boldsymbol{b}-\boldsymbol{\delta}\boldsymbol{b}-(t_{0}+\delta_{0})\boldsymbol{p}_{0}}\right\|_{2}^{2}$
$\qquad\qquad\qquad=\left\|{\left(\bm{A}\bm{D}(\boldsymbol{p}_{0})\hat{\boldsymbol{u}}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})-\boldsymbol{b}-t\boldsymbol{p}_{0}\right)+\bm{A}\bm{D}(\boldsymbol{p}_{0})\boldsymbol{v}-\boldsymbol{\delta}\boldsymbol{b}-\delta_{0}\boldsymbol{p}_{0}}\right\|_{2}^{2}$
$\qquad\qquad\qquad=\left\|{\boldsymbol{d}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})+\bm{A}\bm{D}(\boldsymbol{p}_{0})\boldsymbol{v}-\boldsymbol{\delta}\boldsymbol{b}-\delta_{0}\boldsymbol{p}_{0}}\right\|_{2}^{2},$
From this, we obtain (31) and (32).
Part 2. Set $\boldsymbol{d}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})=\boldsymbol{0}$, $\delta_{0}=t-t_{0}$ with $t\in[0,t_{0}]$ and $\boldsymbol{\delta}\boldsymbol{b}=-(t_{0}-t)\boldsymbol{q}$ in Part 1, then factor out the term $(1-t/t_{0})$ outside the optimization problem to obtain immediately (33) and (34). Note that the case $t_{0}=0$ works because we assume $t\in[0,t_{0}]$.
B.4 Proof of Proposition 5.1
Part 1. First, we invoke Lemma 5.1(ii) with
$\boldsymbol{p}_{0}=\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}),\ \hat{\boldsymbol{u}}(\boldsymbol{p}_{0};t_{0},\boldsymbol{b})=\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\boldsymbol{x}^{s}(t_{0},\boldsymbol{b}),$
and use the optimality conditions $\boldsymbol{d}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b});t_{0},\boldsymbol{b})=\boldsymbol{0}$ to simplify formula (31) to
$\boldsymbol{d}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})$ $;t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q})=$ (69)
$(1-t/t_{0})(\bm{A}\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\hat{\boldsymbol{v}}(t_{0},t,\boldsymbol{b},\boldsymbol{q})+t_{0}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})+\boldsymbol{q})),$
where
$\hat{\boldsymbol{v}}(t_{0},t,\boldsymbol{b},\boldsymbol{q})\in\operatorname*{arg\,min}_{\boldsymbol{v}\in\mathbb{R}^{n}}\left\|{\bm{A}\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\boldsymbol{v}+t_{0}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})+\boldsymbol{q})}\right\|_{2}^{2}$ (70)
subject $\text{subject to}\ \begin{cases}&\boldsymbol{v}_{j}\geqslant-|\boldsymbol{x}^{s}_{j}(t_{0},\boldsymbol{b})|/(1-t/t_{0})\ \text{if $j\in\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$ and $\boldsymbol{x}^{s}_{j}(\boldsymbol{b},t_{0})\neq 0$}\\ &\boldsymbol{v}_{j}\geqslant 0\ \text{if $j\in\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$ and $\boldsymbol{x}^{s}_{j}(\boldsymbol{b},t_{0})=0$,}\\ &\boldsymbol{v}_{\mathcal{E}^{\mathsf{C}}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}=\boldsymbol{0}.\end{cases}$
At $t=t_{0}$, problem (70) reduces to (35) with the identification $\hat{\boldsymbol{v}}(t_{0},t_{0},\boldsymbol{b},\boldsymbol{q})\equiv\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})$.
We will now identify how the solution $\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})$ behaves as we decrease $t_{0}$. There are two potential sources of changes: the set of constraints in (70) and the minimal selection principle via the evolution rule (20) in Proposition 3.1. We turn to these two sources in turn.
First, consider the constraints in problem (70):
[…]
By assumption that $t_{0}>0$ and continuity, there is some $\epsilon_{0}>0$ such that $0\leqslant t_{0}-\epsilon_{0}$, $t\in[t_{0}-\epsilon_{0},t_{0}]$, and $\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})$ remains the solution to problem (70). It will remain the same as $\epsilon_{0}$ decreases until either $\epsilon_{0}=t_{0}$ or there exists $j\in\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$ with $|\boldsymbol{x}^{s}_{j}(t_{0},\boldsymbol{b})|\neq\boldsymbol{0}$ such that $\hat{\boldsymbol{v}}^{s}_{j}(t_{0},\boldsymbol{b},\boldsymbol{q})=-|\boldsymbol{x}^{s}_{j}(t_{0},\boldsymbol{b})|/(1-t/t_{0})$, that is, a constraint becomes satisfied with equality. In the latter case, we can rearrange this expression to obtain
$t=t_{0}\left(1-\frac{|\boldsymbol{x}^{s}_{j}(t_{0},\boldsymbol{b})|}{|\hat{\boldsymbol{v}}^{s}_{j}(t_{0},\boldsymbol{b},\boldsymbol{q})|}\right).$
The first index $j\in\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$ with $\boldsymbol{x}^{s}_{j}(t_{0},\boldsymbol{b})\neq 0$ and $\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})\leqslant-|\boldsymbol{x}^{s}_{j}(t_{0},\boldsymbol{b})|$, if it exists, is the one whose ratio $|\boldsymbol{x}^{s}_{j}(t_{0},\boldsymbol{b})|/|\hat{\boldsymbol{v}}^{s}_{j}(t_{0},\boldsymbol{b},\boldsymbol{q})|$ is minimized. Hence
$T_{-}(t_{0},\boldsymbol{b},\boldsymbol{q},\hat{\boldsymbol{v}}^{s})\coloneqq t_{0}\left(1-\inf_{\begin{subarray}{c}j\in\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\\ \boldsymbol{x}^{s}_{j}(t_{0},\boldsymbol{b})\neq 0\\ \hat{\boldsymbol{v}}_{j}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})\leqslant-|\boldsymbol{x}_{j}^{s}(t_{0},\boldsymbol{b})|\end{subarray}}\frac{|\boldsymbol{x}_{j}^{s}(t_{0},\boldsymbol{b})|}{|\hat{\boldsymbol{v}}^{s}_{j}(t_{0},\boldsymbol{b},\boldsymbol{q})|}\right)$
is the smallest number lesser than $t_{0}$ for which $\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})$ solves problem (70).
Now, consider the descent direction (69), suppose $t\in[\max(0,T_{-}(t_{0},\boldsymbol{b},\boldsymbol{q},\hat{\boldsymbol{v}}^{s})),t_{0}]$ and let
$\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})\coloneqq\bm{A}\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})+t_{0}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})+\boldsymbol{q})$
so as to write
$\boldsymbol{d}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b});t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q})=(1-t/t_{0})\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q}).$
The descent direction depends linearly on $t$, and so the corresponding maximal descent time $\Delta_{*}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b});t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q})$ in Proposition 3.1(i) is inversely proportional to $t$. More precisely:
$\Delta_{*}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b});t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q})=C^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})/(1-t/t_{0}).$ (71)
where
$C^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})\coloneqq$
$\qquad\min_{j\in\{1,\dots,n\}}\left\{\frac{\mathrm{sgn}{\left(\langle\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\bm{A}^{\top}\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q}),\boldsymbol{e}_{j}\rangle\right)}-\langle\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\bm{A}^{\top}\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}),\boldsymbol{e}_{j}\rangle)}{\langle\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\bm{A}^{\top}\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q}),\boldsymbol{e}_{j}\rangle}\right\}.$
Furthermore, the evolution rule from Proposition 3.1(v) yields
$\boldsymbol{d}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})+\Delta(1-t/t_{0})\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q});t,\boldsymbol{b}-(t-t_{0})\boldsymbol{q})=(1-t\Delta)(1-t/t_{0})\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q}).$
We now seek the smallest nonnegative number $T_{+}(t_{0},\boldsymbol{b},\boldsymbol{q})\leqslant t_{0}$ in terms of $\Delta_{*}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b});t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q})$ for which
$1-t\Delta_{*}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b});t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q})=\boldsymbol{0}$
for every $t\in[T_{+}(t_{0},\boldsymbol{b},\boldsymbol{q}),t_{0}]$. Equation (71) gives
$1-t\Delta_{*}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b});t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q})=\boldsymbol{0}$ $\iff 1-tC^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})/(1-t/t_{0})=0$
$\iff 1-t/t_{0}+tC^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})=0$
$\iff t=t_{0}/(1+t_{0}C^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})).$
The critical value is
$T_{+}(t_{0},\boldsymbol{b},\boldsymbol{q})\coloneqq\frac{t_{0}}{1+t_{0}C^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})}.$
Hence letting
$t_{1}\coloneqq\max(T_{-}(t_{0},\boldsymbol{b},\boldsymbol{q},\hat{\boldsymbol{v}}^{s}),T_{+}(t_{0},\boldsymbol{b},\boldsymbol{q})),$
for every $t\in(t_{1},t_{0}]$ there exists some $\Delta\in[0,\Delta_{*}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b});t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q}))$ for which $01-t\Delta=$ 0. Note that $0\leqslant t_{1}<t_{0}$ since $0\leqslant T_{+}(t_{0},\boldsymbol{b},\boldsymbol{q})<t_{0}$ and $T_{-}(t_{0},\boldsymbol{b},\boldsymbol{q},\hat{\boldsymbol{v}}^{s})<t_{0}$. In particular, we have that $t_{1}=0\implies T_{+}(t_{0},\boldsymbol{b},\boldsymbol{q})=0$ and
$T_{+}(t_{0},\boldsymbol{b},\boldsymbol{q})=0\iff C^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})=+\infty\iff\bm{A}^{\top}\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})=\boldsymbol{0}.$
Taken together, we arrive at the following result: For every $t\in(t_{1},t_{0}]$, we have
$\boldsymbol{d}\left(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})+\left(\frac{1}{t}-\frac{1}{t_{0}}\right)\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q});t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q}\right)=\boldsymbol{0}.$
Using the optimality conditions (6) and Lemma 4.1, we conclude
$\boldsymbol{p}^{s}(t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q})=\begin{cases}&\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})+\left(\frac{1}{t}-\frac{1}{t_{0}}\right)\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})\quad\text{if $t_{1}>0$},\\ &\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})\quad\text{otherwise,}\end{cases}$
is the solution to (dBPDN) at hyperparameter $t$ and data $\boldsymbol{b}-(t_{0}-t)\boldsymbol{q}$ on $[t_{1},t_{0}]$. Furthermore,
$t\boldsymbol{p}^{s}(t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q})$ $=\left(\frac{t}{t_{0}}\right)t_{0}\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})+\left(1-\frac{t}{t_{0}}\right)\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})$
$=t_{0}\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})-\left(1-\frac{t}{t_{0}}\right)t_{0}\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})$
$\qquad\qquad+\left(1-\frac{t}{t_{0}}\right)(\bm{A}\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))+t_{0}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})+\boldsymbol{q}))$
$=t_{0}\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})+\left(1-\frac{t}{t_{0}}\right)\left(\bm{A}\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})+t_{0}\boldsymbol{q}\right)$
$=\bm{A}\left[\boldsymbol{x}^{s}(t_{0},\boldsymbol{b})+\left(1-\frac{t}{t_{0}}\right)\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})\right]-\boldsymbol{b}-(t_{0}-t)\boldsymbol{q},$
and so we deduce
$\boldsymbol{x}^{s}(t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q})=\boldsymbol{x}^{s}(t_{0},\boldsymbol{b})+\left(1-\frac{t}{t_{0}}\right)\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})$
is the primal solution to (BPDN) at hyperparameter $t$ and data $\boldsymbol{b}-(t_{0}-t)\boldsymbol{q}$ on $[t_{1},t_{0}]$.
Part 2. For the rest of the proof, we will write, for brevity,
$\boldsymbol{b}_{t}\equiv\boldsymbol{b}-(t-t_{0})\boldsymbol{q},\quad\boldsymbol{\xi}^{s}_{0}\equiv\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q}),$
$\boldsymbol{p}_{t}\equiv\boldsymbol{p}^{s}(t,\boldsymbol{b}-(t-t_{0})\boldsymbol{q})\quad\text{and}\quad\boldsymbol{x}_{t}=\boldsymbol{x}^{s}(t,\boldsymbol{b}-(t-t_{0})\boldsymbol{q}).$
First, we use Proposition 3.1(ii) with $\Delta=1/t$ to evaluate $V(\boldsymbol{p}_{t};t,\boldsymbol{b}_{t})$:
$V(\boldsymbol{p}_{t};t,\boldsymbol{b}_{t})=V(\boldsymbol{p}_{t_{0}};t,\boldsymbol{b}_{t})-\frac{1}{2t}\left(1-\frac{t}{t_{0}}\right)^{2}\left\|{\boldsymbol{\xi}_{0}^{s}}\right\|_{2}^{2}.$
We can develop the first term on the right hand side to get
$V(\boldsymbol{p}_{t_{0}};t,\boldsymbol{b}_{t})$ $=\frac{t}{2}\left\|{\boldsymbol{p}_{t_{0}}}\right\|_{2}^{2}+\langle\boldsymbol{p}_{t_{0}},\boldsymbol{b}_{t}\rangle$
$=\frac{t_{0}}{2}\left\|{\boldsymbol{p}_{t_{0}}}\right\|_{2}^{2}+\langle\boldsymbol{p}_{t_{0}},\boldsymbol{b}_{t_{0}}\rangle+\left(\frac{t-t_{0}}{2}\right)\left\|{\boldsymbol{p}_{t_{0}}}\right\|_{2}^{2}+(t-t_{0})\langle\boldsymbol{p}_{t_{0}},\boldsymbol{q}\rangle$
$=V(\boldsymbol{p}_{t_{0}};t_{0},\boldsymbol{b}_{t_{0}})+\left(\frac{t-t_{0}}{2}\right)\left\|{\boldsymbol{p}_{t_{0}}}\right\|_{2}^{2}+(t-t_{0})\langle\boldsymbol{p}_{t_{0}},\boldsymbol{q}\rangle.$
Combining the two equations above yields
$V(\boldsymbol{p}_{t};t,\boldsymbol{b}_{t})-V(\boldsymbol{p}_{t_{0}};t_{0},\boldsymbol{b}_{t_{0}})=-\frac{1}{2t}\left(1-\frac{t}{t_{0}}\right)^{2}\left\|{\boldsymbol{\xi}_{0}^{s}}\right\|_{2}^{2}-(t_{0}-t)\left(\frac{1}{2}\left\|{\boldsymbol{p}_{t_{0}}}\right\|_{2}^{2}+\langle\boldsymbol{p}_{t_{0}},\boldsymbol{q}\rangle\right).$ (72)
Now, let $\boldsymbol{q}=-\bm{A}\bm{D}(\boldsymbol{p}_{t_{0}})\boldsymbol{w}/t_{0}$ for some nonnegative $\boldsymbol{w}\in\mathbb{R}^{n}$. We have
$\langle\boldsymbol{p}_{t_{0}},\boldsymbol{q}\rangle=\langle-\bm{D}(\boldsymbol{p}_{t_{0}})\bm{A}^{\top}\boldsymbol{p}_{t_{0}},\boldsymbol{w}\rangle/t_{0}\geqslant 0$
because $\boldsymbol{0}\leqslant-\bm{D}(\boldsymbol{p}_{t_{0}})\bm{A}^{\top}\boldsymbol{p}_{t_{0}}\leqslant\boldsymbol{1}$. The inequality for the dual objective function $V$ in (38) follows immediately, and the inequality is strict if $\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b},\boldsymbol{q})\neq\boldsymbol{0}$.
Next, we use the optimality of the dual solutions $\boldsymbol{p}_{t_{0}}$ and $\boldsymbol{p}_{t}$ to obtain the two inequalities
$\frac{t_{0}}{2}\left\|{\boldsymbol{p}_{t_{0}}}\right\|_{2}^{2}+\langle\boldsymbol{p}_{t_{0}},\boldsymbol{b}_{t_{0}}\rangle$ $\leqslant\frac{t_{0}}{2}\left\|{\boldsymbol{p}_{t}}\right\|_{2}^{2}+\langle\boldsymbol{p}_{t},\boldsymbol{b}_{t_{0}}\rangle,$ (73)
$\frac{t}{2}\left\|{\boldsymbol{p}_{t}}\right\|_{2}^{2}+\langle\boldsymbol{p}_{t},\boldsymbol{b}_{t}\rangle$ $\leqslant\frac{t}{2}\left\|{\boldsymbol{p}_{0}}\right\|_{2}^{2}+\langle\boldsymbol{p}_{0},\boldsymbol{b}_{t}\rangle.$
Writing
$\langle\boldsymbol{p}_{t},\boldsymbol{b}_{t}\rangle$ $=\langle\boldsymbol{p}_{t},\boldsymbol{b}_{t_{0}}\rangle-(t_{0}-t)\langle\boldsymbol{p}_{t},\boldsymbol{q}\rangle,$
$\langle\boldsymbol{p}_{0},\boldsymbol{b}_{t}\rangle$ $=\langle\boldsymbol{p}_{0},\boldsymbol{b}_{t_{0}}\rangle-(t_{0}-t)\langle\boldsymbol{p}_{0},\boldsymbol{q}\rangle,$
we can rearrange the second inequality as
$\langle\boldsymbol{p}_{t},\boldsymbol{b}_{t_{0}}\rangle$ $\leqslant\frac{t}{2}\left\|{\boldsymbol{p}_{t_{0}}}\right\|_{2}^{2}-\frac{t}{2}\left\|{\boldsymbol{p}_{t}}\right\|_{2}^{2}+\langle\boldsymbol{p}_{t_{0}},\boldsymbol{b}_{t}\rangle+(t_{0}-t)\langle\boldsymbol{p}_{t},\boldsymbol{q}\rangle$
$=\frac{t}{2}\left\|{\boldsymbol{p}_{t_{0}}}\right\|_{2}^{2}-\frac{t}{2}\left\|{\boldsymbol{p}_{t}}\right\|_{2}^{2}+\langle\boldsymbol{p}_{t_{0}},\boldsymbol{b}_{t_{0}}\rangle+(t_{0}-t)\langle\boldsymbol{p}_{t}-\boldsymbol{p}_{t_{0}},\boldsymbol{q}\rangle.$
Using this in the first inequality in (73) yields
$\frac{t_{0}}{2}\left\|{\boldsymbol{p}_{t_{0}}}\right\|_{2}^{2}+\langle\boldsymbol{p}_{t_{0}},\boldsymbol{b}_{t_{0}}\rangle\leqslant\frac{t_{0}}{2}\left\|{\boldsymbol{p}_{t}}\right\|_{2}^{2}+\frac{t}{2}\left\|{\boldsymbol{p}_{t_{0}}}\right\|_{2}^{2}-\frac{t}{2}\left\|{\boldsymbol{p}_{t}}\right\|_{2}^{2}+\langle\boldsymbol{p}_{t_{0}},\boldsymbol{b}_{t_{0}}\rangle+(t_{0}-t)\langle\boldsymbol{p}_{t}-\boldsymbol{p}_{t_{0}},\boldsymbol{q}\rangle.$
Rearranging and factoring out the positive term $t_{0}-t$, this simplifies to
$\frac{1}{2}\left\|{\boldsymbol{p}_{t_{0}}}\right\|_{2}^{2}\leqslant\frac{1}{2}\left\|{\boldsymbol{p}_{t}}\right\|_{2}^{2}-\langle\bm{D}(\boldsymbol{p}_{t_{0}})\bm{A}^{\top}(\boldsymbol{p}_{t}-\boldsymbol{p}_{t_{0}}),\boldsymbol{w}\rangle/t_{0}.$ (74)
If $\boldsymbol{p}_{t}=\boldsymbol{p}_{t_{0}}$, which is equivalent to $\boldsymbol{\xi}^{s}_{0}=\boldsymbol{0}$, then the second term on the right hand side vanishes and we obtain the second inequality in (38). If not, we can further write (74) as
$\frac{1}{2}\left\|{\boldsymbol{p}_{t_{0}}}\right\|_{2}^{2}\leqslant\frac{1}{2}\left\|{\boldsymbol{p}_{t}}\right\|_{2}^{2}-\left(\frac{1}{t}-\frac{1}{t_{0}}\right)\langle\bm{D}(\boldsymbol{p}_{t_{0}})\bm{A}^{\top}\boldsymbol{\xi}^{s}_{0},\boldsymbol{w}\rangle/t_{0}.$
The optimality conditions of the NNLS problem (35) yield $\bm{D}(\boldsymbol{p}_{t_{0}})\bm{A}^{\top}\boldsymbol{\xi}^{s}_{0}\geqslant\boldsymbol{0}$, and consequently $\langle\bm{D}(\boldsymbol{p}_{t_{0}})\bm{A}^{\top}\boldsymbol{\xi}^{s}_{0},\boldsymbol{w}\rangle/t_{0}\geqslant 0$ by assumption. We deduce
$\left\|{\boldsymbol{p}_{t_{0}}}\right\|_{2}\leqslant\left\|{\boldsymbol{p}_{t}}\right\|_{2}$ (75)
If $\boldsymbol{\xi}^{s}_{0}\neq\boldsymbol{0}$, then $\boldsymbol{p}_{t}\neq\boldsymbol{p}_{t_{0}}$ and $t_{1}>0$, so the inequalities in (73) become strict by uniqueness of solutions to (dBPDN). The strict inequalities propagate to (75), proving the result.
B.5 Proof of Proposition 6.1
We first show the equicorrelation set enforces a consistency property.
Lemma B.1.
Let $t_{0}>0,\boldsymbol{p}_{0}\in\mathrm{dom~{}}\partial_{\boldsymbol{p}}V(\cdot;t_{0},\boldsymbol{b})$, and suppose the equicorrelation set $\mathcal{E}(\boldsymbol{p}_{0})$ is non-empty. Consider the least-squares problem with unknown $\boldsymbol{v}$ and offset $\boldsymbol{w}\in\mathbb{R}^{|\mathcal{E}(\boldsymbol{p}_{0})|}$,
$\min_{\boldsymbol{v}\in\mathbb{R}^{|\mathcal{E}(\boldsymbol{p}_{0})|}}\left\|{\bm{A}_{\mathcal{E}(\boldsymbol{p}_{0})}\bm{D}_{\mathcal{E}(\boldsymbol{p}_{0})}(\boldsymbol{v}-\boldsymbol{w})+t_{0}\boldsymbol{p}_{0}}\right\|_{2}^{2}.$ (76)
Letting
$\mathcal{M}(\boldsymbol{p}_{0})\in\sup_{|\mathcal{M}|}\left\{\text{$\mathcal{M}\subset\mathcal{E}(\boldsymbol{p}_{0})$ : s.t. $\left\{\bm{A}\boldsymbol{e}_{j}\right\}_{j\in\mathcal{M}}$ are linearly independent}\right\},$ (77)
the least-squares problem (76) admits the solution
$\hat{\boldsymbol{v}}\coloneqq\begin{cases}\hat{\boldsymbol{v}}_{\mathcal{M}(\boldsymbol{p}_{0})}&=\boldsymbol{w}_{\mathcal{M}(\boldsymbol{p}_{0})}+t_{0}\left(\bm{A}^{\top}_{\mathcal{M}(\boldsymbol{p}_{0})}\bm{A}_{\mathcal{M}(\boldsymbol{p}_{0})}\right)^{-1}\boldsymbol{1}_{\mathcal{M}(\boldsymbol{p}_{0})},\\ \hat{\boldsymbol{v}}_{\mathcal{E}{(\boldsymbol{p})}\setminus\mathcal{M}(\boldsymbol{p}_{0})}&=\boldsymbol{w}_{\mathcal{E}(\boldsymbol{p})\setminus\mathcal{M}(\boldsymbol{p}_{0})}.\end{cases}$ (78)
Moreover, the residual vector $\boldsymbol{\xi}$ of the least-squares problem (76) is independent of $\boldsymbol{w}$:
$\boldsymbol{\xi}=t_{0}\bm{A}_{\mathcal{M}(\boldsymbol{p}_{0})}\bm{D}_{\mathcal{M}(\boldsymbol{p}_{0})}\left(\bm{A}^{\top}_{\mathcal{M}(\boldsymbol{p}_{0})}\bm{A}_{\mathcal{M}(\boldsymbol{p}_{0})}\right)^{-1}\boldsymbol{1}_{\mathcal{M}(\boldsymbol{p}_{0})}.$ (79)
Proof.
The least-squares problem (76) has at least one solution and its solution(s) satisfy the normal equations
$\bm{A}^{\top}_{\mathcal{E}(\boldsymbol{p}_{0})}\bm{A}_{\mathcal{E}(\boldsymbol{p}_{0})}(\boldsymbol{v}-\boldsymbol{w})+t_{0}\bm{D}_{\mathcal{E}(\boldsymbol{p}_{0})}\bm{A}^{\top}_{\mathcal{E}(\boldsymbol{p}_{0})}\boldsymbol{p}_{0}=\boldsymbol{0}.$
(This follows from Theorem A.3 with $f_{1}=\boldsymbol{0}$ and $f_{2}=\frac{1}{2}\left\|{\cdot}\right\|_{2}^{2}$ and the generalized Fermat’s rule A.1.) The equicorrelation property $-\bm{D}_{\mathcal{E}(\boldsymbol{p}_{0})}\bm{A}^{\top}_{\mathcal{E}(\boldsymbol{p}_{0})}\boldsymbol{p}_{0}=\boldsymbol{1}_{\mathcal{E}(\boldsymbol{p}_{0})}$ simplifies the above to
$\bm{A}^{\top}_{\mathcal{E}(\boldsymbol{p}_{0})}\bm{A}_{\mathcal{E}(\boldsymbol{p}_{0})}(\boldsymbol{v}-\boldsymbol{w})=t_{0}\boldsymbol{1}_{\mathcal{E}(\boldsymbol{p}_{0})}.$ (80)
Now, consider the set in (77) on the right. It is non-empty and so the supremum is attained because the equicorrelation set is finite. Hence $\mathcal{M}(\boldsymbol{p}_{0})$ is well-defined. We now show the vector $\hat{\boldsymbol{v}}$ defined in (78) and formed from $\mathcal{M}(\boldsymbol{p}_{0})$ solves the normal equations (80). First, suppose $\mathcal{M}(\boldsymbol{p}_{0})=\mathcal{E}(\boldsymbol{p}_{0})$. Since $\mathrm{rank}({\bm{A}})=m$ by assumption, we have $|\mathcal{E}(\boldsymbol{p}_{0})|\leqslant m$ and so $\bm{A}_{\mathcal{E}(\boldsymbol{p}_{0})}$ has full column rank. Hence $\bm{A}^{\top}_{\mathcal{E}(\boldsymbol{p}_{0})}\bm{A}_{\mathcal{E}(\boldsymbol{p}_{0})}$ is invertible (see [8, Sections 5.12-5.13, pages 187-190], and note that invertibility is equivalent to $\bm{A}_{\mathcal{E}(\boldsymbol{p}_{0})}$ having full column rank, itself equivalent to injectivity of $\bm{A}_{\mathcal{E}(\boldsymbol{p}_{0})}$) and a simple calculation shows the claim.
Next, suppose $(\mathcal{E}(\boldsymbol{p}_{0})\setminus\mathcal{M}(\boldsymbol{p}_{0}))\neq\emptyset$ and let $i\in\mathcal{E}(\boldsymbol{p}_{0})\setminus\mathcal{M}(\boldsymbol{p}_{0})$. There are coefficients $\{c_{ij}\}_{j\in\mathcal{M}(\boldsymbol{p}_{0})}$, not all zero, such that
$\bm{A}\boldsymbol{e}_{i}=\sum_{j\in\mathcal{M}(\boldsymbol{p}_{0})}c_{ij}\bm{A}\boldsymbol{e}_{j}.$
Multiply the equation above by the matrix of signs $\bm{D}(\boldsymbol{p}_{0})$, take the scalar product on both sides with respect to $-\boldsymbol{p}_{0}$ and rearrange to find
$\langle-\bm{D}(\boldsymbol{p}_{0})\bm{A}^{\top}\boldsymbol{p}_{0},\boldsymbol{e}_{i}\rangle=\sum_{j\in\mathcal{M}(\boldsymbol{p}_{0})}c_{ij}\langle-\bm{D}(\boldsymbol{p}_{0})\bm{A}^{\top}\boldsymbol{p},\boldsymbol{e}_{j}\rangle.$
However, the equicorrelation property implies
$1=\sum_{j\in\mathcal{M}(\boldsymbol{p}_{0})}c_{ij}.$ (81)
That is, each column vector $\bm{A}\boldsymbol{e}_{i}$ with $i\in\mathcal{E}(\boldsymbol{p}_{0})\setminus\mathcal{M}(\boldsymbol{p}_{0})$ satisfies the constraint (81). Suppose, without loss of generality, the $i^{\text{th}}$ row of the augmented matrix $(\bm{A}^{\top}_{\mathcal{E}(\boldsymbol{p}_{0})}\bm{A}_{\mathcal{E}(\boldsymbol{p}_{0})}|t_{0}\boldsymbol{1}_{\mathcal{E}(\boldsymbol{p}_{0})})$ is formed using the column vector $\bm{A}\boldsymbol{e}_{i}$ with $i\in\mathcal{E}(\boldsymbol{p}_{0})\setminus\mathcal{M}(\boldsymbol{p}_{0})$. The $|\mathcal{E}(\boldsymbol{p}_{0})|+1$ coefficients of this row are the scalar products $\langle\bm{A}\boldsymbol{e}_{i},\bm{A}\boldsymbol{e}_{k}\rangle$ and $1$. The linear dependence of $\bm{A}\boldsymbol{e}_{i}$ on the column vectors $\{\bm{A}\boldsymbol{e}_{j}\}_{j\in\mathcal{M}(\boldsymbol{p}_{0})}$ implies
$t_{0}\langle\bm{A}\boldsymbol{e}_{i},\bm{A}\boldsymbol{e}_{k}\rangle-t_{0}\sum_{j\in\mathcal{M}(\boldsymbol{p}_{0})}c_{ij}\langle\bm{A}\boldsymbol{e}_{j},\bm{A}\boldsymbol{e}_{k}\rangle=0\ \text{for every $k\in\mathcal{E}(\boldsymbol{p}_{0})$.}$
In particular, we can eliminate the first $|\mathcal{E}(\boldsymbol{p}_{0})|$ coefficients of the $i^{\text{th}}$ row by subtracting multiples $t_{0}c_{ij}$ of the augmented matrix’s rows formed by the column vectors $\bm{A}\boldsymbol{e}_{j}$ indexed by $\mathcal{M}(\boldsymbol{p}_{0})$. Crucially, the last element of the row is also eliminated because $t_{0}(1-\sum_{j\in\mathcal{M}(\boldsymbol{p}_{0})}c_{ij})=0$ thanks to (81). Since $i\in\mathcal{E}(\boldsymbol{p}_{0})\setminus\mathcal{M}(\boldsymbol{p}_{0})$ was arbitrary, all rows of the augmented matrix $(\bm{A}^{\top}_{\mathcal{E}(\boldsymbol{p}_{0})}\bm{A}_{\mathcal{E}(\boldsymbol{p}_{0})}|t_{0}\boldsymbol{1}_{\mathcal{E}(\boldsymbol{p}_{0})})$ formed by the column vectors $\{\bm{A}\boldsymbol{e}_{i}\}_{i\in\mathcal{E}(\boldsymbol{p})\setminus\mathcal{M}(\boldsymbol{p}_{0})}$ can be eliminated consistently. It therefore suffices to compute a solution to the reduced linear system $\bm{A}^{\top}_{\mathcal{\boldsymbol{p}}}\bm{A}_{\mathcal{\boldsymbol{p}}}\boldsymbol{v}=t_{0}\boldsymbol{1}_{\mathcal{M}}$ and add the offset $\boldsymbol{w}$ to get the solution to the linear system (80). ∎
Proof of Proposition 6.1(i). Let $\boldsymbol{w}\in\mathbb{R}^{n}$ and $\boldsymbol{q}=-\bm{A}\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\boldsymbol{w}/t_{0}$. Consider the NNLS problem (35) without the constraints “$\boldsymbol{v}_{j}\geqslant 0\ \text{if $j\in\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$ and $\boldsymbol{x}_{j}^{s}(t_{0},\boldsymbol{b})=0$}$”:
$\min_{\begin{subarray}{c}\boldsymbol{v}_{\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}\in\mathbb{R}^{|\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))|}\\ \boldsymbol{v}_{\mathcal{E}^{\mathsf{C}}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}=\boldsymbol{w}_{\mathcal{E}^{\mathsf{C}}(\boldsymbol{p}^{s})(t_{0},\boldsymbol{b})}\end{subarray}}\left\|{\bm{A}_{\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}\bm{D}_{\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}(\boldsymbol{v}_{\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}-\boldsymbol{w}_{\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))})+t_{0}\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})}\right\|_{2}^{2}.$ (82)
Using Lemma B.1, the set of solutions to (82) contains $\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b},\boldsymbol{w})=\hat{\boldsymbol{v}}^{\text{LSQ}}(t_{0},\boldsymbol{b})+\boldsymbol{w}$, with $\hat{\boldsymbol{v}}^{\text{LSQ}}(t_{0},\boldsymbol{b})$ given by (39). If $\boldsymbol{w}$ also satisfies (42), then $\hat{\boldsymbol{v}}^{s}_{j}(t_{0},\boldsymbol{b},\boldsymbol{w})\geqslant 0$ for every $j\in\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$ with $\boldsymbol{x}_{j}^{s}(t_{0},\boldsymbol{b})=0$. Hence $\hat{\boldsymbol{v}}^{s}$ satisfies the constraints of the NNLS problem (35) and is one of its solutions. The explicit formula for $\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b})$ on the first line of (43) follows by substituting the solution $\hat{\boldsymbol{v}}^{s}$ into $\bm{A}\bm{D}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))(\boldsymbol{v}-\boldsymbol{w})+t_{0}\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})$. The equation $\bm{A}^{\top}_{\mathcal{E}{(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}}\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b})=\boldsymbol{0}$ follows from the first-order optimality conditions of (82), and the breakpoint $T_{+}$ is independent of $\boldsymbol{w}$ because $\boldsymbol{\xi}^{s}$ itself is independent of $\boldsymbol{w}$.
Consider now the breakpoint $T_{-}(t_{0},\boldsymbol{b},\boldsymbol{w},\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b},\boldsymbol{w}))$ in (36). Since $\boldsymbol{w}$ satisfies (42),
$\inf_{\begin{subarray}{c}j\in\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\\ \boldsymbol{x}^{s}(t_{0},\boldsymbol{b})\neq 0\\ \hat{\boldsymbol{v}}^{s}_{j}(t_{0},\boldsymbol{b},\boldsymbol{w})\leqslant-|\boldsymbol{x}_{j}^{s}(t_{0},\boldsymbol{b})|\end{subarray}}\frac{|\boldsymbol{x}_{j}^{s}(t_{0},\boldsymbol{b})|}{|\hat{\boldsymbol{v}}^{s}_{j}(t_{0},\boldsymbol{b},\boldsymbol{w})|}\geqslant 1,$
[…]
$T_{-}(t_{0},\boldsymbol{b},\boldsymbol{w},\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b},\boldsymbol{w}))\leqslant 0\leqslant T_{+}(t_{0},\boldsymbol{b}).$
Proof of Proposition 6.1(ii). Let $k\in\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$. By Part 1, $T_{-}(t_{0},\boldsymbol{b},\boldsymbol{w},\hat{\boldsymbol{v}}^{s}(t_{0},\boldsymbol{b},\boldsymbol{w}))\leqslant T_{+}(t_{0},\boldsymbol{b})$, so we can invoke Proposition 5.1 and formula (37) with $t\in(T_{+}(t_{0},\boldsymbol{b}),t_{0}]$ to write:
$\left\langle-\bm{A}^{\top}\boldsymbol{p}^{s}(t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q}),\boldsymbol{e}_{k}\right\rangle$ $=\left\langle-\bm{A}^{\top}\left(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})-\left(\frac{1}{t}-\frac{1}{t_{0}}\right)\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b})\right),\boldsymbol{e}_{k}\right\rangle$
$=\left\langle-\bm{A}^{\top}\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}),\boldsymbol{e}_{k}\right\rangle-\left(\frac{1}{t}-\frac{1}{t_{0}}\right)\left\langle\bm{A}^{\top}\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b}),\boldsymbol{e}_{k}\right\rangle$
$=\left\langle-\bm{A}^{\top}\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}),\boldsymbol{e}_{k}\right\rangle$
$=\mathrm{sgn}(\left\langle-\bm{A}^{\top}\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}),\boldsymbol{e}_{k}\right\rangle),$
where the penultimate line follows from $\bm{A}^{\top}_{\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}\boldsymbol{\xi}^{s}(t_{0},\boldsymbol{b})=\boldsymbol{0}$. Hence we find
$\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\subset\mathcal{E}\left(\boldsymbol{p}^{s}(t,\boldsymbol{b}-(t_{0}-t)\boldsymbol{q})\right)$
Combined with Proposition 3.1(iv), we obtain the set equality (44). If $T_{+}(t_{0},\boldsymbol{b})>0$, then by continuity we have the inclusion
$\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))\subset\mathcal{E}(\boldsymbol{p}^{s}(T_{+}(t_{0},\boldsymbol{b}),\boldsymbol{b}-(t_{0}-T_{+}(t_{0},\boldsymbol{b}))\boldsymbol{q})).$
At $t=T_{+}(t_{0},\boldsymbol{b})$, a new element $j\in\mathcal{E}^{\mathsf{C}}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$ is added by Propositions 3.1 and 5.1, and so the strict inclusion (45) follows.
Proof of Proposition 6.1(iii). For the final part of the proof, we will write, for brevity,
$\boldsymbol{p}_{+}$ $\equiv\boldsymbol{p}^{s}(T_{+}(t_{0},\boldsymbol{b}),\boldsymbol{b}-(t_{0}-T_{+}(t_{0},\boldsymbol{b}))\boldsymbol{q}),$
$\bm{D}_{+}$ $\equiv\bm{D}(\boldsymbol{p}^{s}(T_{+}(t_{0},\boldsymbol{b}),\boldsymbol{b}-(t_{0}-T_{+}(t_{0},\boldsymbol{b}))\boldsymbol{q})).$
Consider the set difference $\mathcal{E}\left(\boldsymbol{p}_{+}\right)\setminus\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$, which is non-empty, and let $j\in\mathcal{E}\left(\boldsymbol{p}_{+}\right)\setminus\mathcal{E}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$. Assume for a contradiction that $\bm{A}\bm{D}_{+}\boldsymbol{e}_{j}$ depends linearly on the vectors $\{\bm{A}\bm{D}_{+}\boldsymbol{e}_{k}\}_{k\in\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}$. Then there exist numbers $\{c_{k}\}_{k\in\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}$, not all zero, such that
$\sum_{k\in\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}c_{k}(-\bm{A}\bm{D}_{+}\boldsymbol{e}_{k})-\bm{A}\bm{D}_{+}\boldsymbol{e}_{j}=\boldsymbol{0}.$ (83)
Take the scalar product with respect to $\boldsymbol{p}_{+}$ in (83) and use the set inclusion (45) to find
$\sum_{k\in\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}c_{k}\langle-\bm{D}_{+}\bm{A}^{\top}\boldsymbol{p}_{+},\boldsymbol{e}_{k}\rangle$ $+\langle-\bm{D}_{+}\bm{A}^{\top}\boldsymbol{p}_{+},\boldsymbol{e}_{j}\rangle=\sum_{k\in\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}c_{k}+1=0$
$\iff\sum_{k\in\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}c_{k}=-1.$
Next, note that $|\langle-\bm{A}^{\top}\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}),\boldsymbol{e}_{j}\rangle|<1$ because $j\notin\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))$. Hence taking the scalar product with respect to $\boldsymbol{p}^{s}(t_{0},\boldsymbol{b})$ in (83) and using, again, the set inclusion (45), we find
$\sum_{k\in\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}c_{k}\langle-\bm{D}_{+}\bm{A}^{\top}\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}),\boldsymbol{e}_{k}\rangle$ $+\langle-\bm{D}_{+}\bm{A}^{\top}\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}),\boldsymbol{e}_{j}\rangle$
$=\sum_{k\in\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}c_{k}+\langle-\bm{D}_{+}\bm{A}^{\top}\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}),\boldsymbol{e}_{j}\rangle=0$
$\iff\sum_{k\in\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}c_{k}=-|\langle\bm{A}^{\top}\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}),\boldsymbol{e}_{j}\rangle|$
$\implies\sum_{k\in\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}c_{k}>-1.$
This gives a contradiction, and the column vector $\bm{A}\boldsymbol{e}_{j}$ is therefore linearly independent from the column vectors $\{\bm{A}\boldsymbol{e}_{k}\}_{k\in\mathcal{M}(\boldsymbol{p}^{s}(t_{0},\boldsymbol{b}))}$.